### v1

# The finite groups with three automorphism orbits

Abstract.
A complete classification is given of finite groups whose elements are partitioned into three orbits by the automorphism groups, solving the long-standing classification problem initiated by G. Higman in 1963. As a consequence, a classification is obtained for finite permutation groups of rank $3$ which are holomorphs of groups.
Key words and phrases:
automorphism group; automorphism orbit; exterior squre
1. Introduction
For a group $N$, the natural action of the automorphism group ${\mathrm{Aut}}(N)$ on $N$ partitions the elements of $N$ into orbits, called the automorphism orbits of $N$. We remark that an automorphism orbit is also called a fusion class. The study of the relations between groups and their automorphism orbits was initiated by G. Higman [16] in 1963 when he studied finite $2$-groups whose involutions form an automorphism orbit, and since then automorphism orbits of groups have received considerable attention of group theorists.
For convenience, a group $N$ is called $r$-orbit if ${\mathrm{Aut}}(N)$ partitions elements of $N$ into $r$ orbits. Since elements in the same automorphism orbit have the same order, the identity group is the only $1$-orbit group, and elementary abelian $p$-groups with $p$ prime are the only $2$-orbit groups. The class of finite $3$-orbit groups contains some interesting families of groups, and the problem of classifying such groups has received considerable attention in the literature. The main purpose of this paper is to classify all finite $3$-orbit groups.
Obviously, a $3$-orbit group has order divisible by at most two primes, and so it is either a $p$-group or a $\{p,q\}$-group, where $p,q$ are distinct primes. It is not hard to see that finite $3$-orbit abelian groups are homocyclic group with exponent $p^{2}$ with $p$ prime. Finite $\{p,q\}$-groups that are $3$-orbit groups have been classified in [10, 12, 20, 24] (also see Example 2.1), which are proved to be certain Frobenius groups. Recently, finite $2$-groups that are $3$-orbit groups are classified independently by Bors and Glasby [4], and the authors [22], with two typical examples being Sylow $2$-subgroups of unitary groups ${\mathrm{SU}}(3,2^{f})$ and Suzuki groups ${\mathrm{Sz}}(2^{2f+1})$ for $f\geqslant 1$ (see Section 2.1). The problem of classifying $3$-orbit groups is therefore reduced to the $p$-groups with odd prime $p$.
As comparison, we would like to mention some interesting researches about automorphism orbits. Studies of $r$-orbit groups for small values of $r>3$ mainly focus on non-solvable groups. Finite simple $r$-orbit groups for $r\leqslant 100$ are determined in [17], and non-solvable $r$-orbit groups for $r\in\{4,5,6\}$ are classified (see [1, 9, 20]). Finite groups with some special automorphism orbits also receive a lot of attention. Shult proved that finite $p$-group for odd prime $p$ such that elements of order $p$ form an automorphism orbit is abelian [27]. In 1976, Gross studied finite $2$-groups whose involutions form an automorphism orbits [14], and the authors complete the classification of such groups recently [22]. Influenced by Gross’ work, finite groups whose elements of the same order lie in the same automorphism orbit, so-called AT-groups, are studied by Zhang [32]. Praeger and the first-named author [21] give a description of finite groups such that elements of the same order are fused or inverse-fused.
Recall that a non-abelian $p$-group $N$ is said to be a special $p$-group if ${\mathbf{Z}}(N)=\Phi(N)=N^{\prime}$ is elementary abelian. In this paper, for a prime power $q=p^{f}$, we denote by $q^{m+n}$ a non-abelian special $p$-group $N$ with $|\Phi(N)|=q^{m}$ and $|N/\Phi(N)|=q^{n}$. We remark that finite non-abelian $3$-orbit $p$-groups are special $p$-groups since it has a unique non-trivial proper characteristic subgroup. For odd prime $p$, finite non-abelian $3$-orbit $p$-groups are special $p$-groups of exponent $p$. It is known that the $n$-generator free group in the variety of groups having exponent $p$ and nilpotency class $2$ is finite, which is denoted by ${\mathcal{H}}_{n,p}$ in this paper (see Definition 2.9). Hence, for odd prime $p$, each non-abelian $n$-generator $3$-orbit $p$-group is isomorphic to a quotient group ${\mathcal{H}}_{n,p}/W$ for some $W<{\mathbf{Z}}({\mathcal{H}}_{n,p})$. To classify $3$-orbit groups, we have to determine all “desired” subgroups $W$. Denote by $\Lambda^{2}_{{\mathrm{F}}}(V)$ the exterior square of vector space $V$ over ${\mathrm{F}}$. We remark that ${\mathbf{Z}}({\mathcal{H}}_{n,p})$ can be “viewed” as the exterior square of ${\mathcal{H}}_{n,p}/{\mathbf{Z}}({\mathcal{H}}_{n,p})$ in some sense (see Lemma 2.10 for more details). The following theorem on representations of finite transitive linear groups plays an important role in the classification.
Theorem A.
Let $V={\mathbb{F}}_{p}^{n}$ for odd prime $p$, $G\leqslant{\mathrm{GL}}(V)$ be non-solvable, and let $W<\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$. If $G$ acts transitively on non-zero vectors of both $V$ and $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$, then one of the following statements holds.
  1. (i)
    $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W=1$.
  2. (ii)
    $G_{0}:=G^{(\infty)}\cong{\mathrm{SL}}(3,p^{n/3})$ and $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W=\dim V=n$.
  3. (iii)
    $G_{0}:=G^{(\infty)}\cong{\mathrm{Sp}}(2m,p^{n/(2m)})$ and $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ is a trivial ${\mathbb{F}}_{p}G_{0}$-module with dimension no more than $n/2m$.
Now we present our main theorem on the classification of finite $3$-orbit groups.
Theorem B.
A finite group is a $3$-orbit group if and only if it is one of the groups listed in Table 1, where $p,q$ are distinct primes, and $m,n$ are positive integers.
Table 1. Finite $3$-orbit groups
$N$ ${\mathrm{Aut}}(N)$ Conditions Ref
(1) ${\mathbb{Z}}_{p}^{n(q-1)}{:}{\mathbb{Z}}_{q}$ ${\mathrm{A}\Gamma\mathrm{L}}(n,p^{q-1})$ Example 2.1 2.2
(2) ${\mathbb{Z}}_{p^{2}}^{n}$ ${\mathrm{GL}}(n,{\mathbb{Z}}/p^{2}{\mathbb{Z}})\cong p^{n^{2}}.{\mathrm{GL}}(n,p)$ 2.3
(3) $A_{2}(n,\theta)\cong 2^{n+n}$ $2^{n^{2}}{:}{\Gamma\mathrm{L}}(1,2^{n})$ $|\theta|\neq 1$ is odd 2.6
(4) ${\mathrm{SU}}(3,2^{n})_{2}\cong 2^{n+2n}$ $2^{2n^{2}}{:}{\Gamma\mathrm{L}}(1,2^{2n})$ 2.6
(5) $P(\epsilon)\cong 2^{3+6}$ $2^{18}{:}({\mathbb{Z}}_{7}{:}{\mathbb{Z}}_{9})$ 2.6
(6) $A_{p}(n,\theta)\cong p^{n+n}$ $p^{n^{2}}{:}{\Gamma\mathrm{L}}(3,p^{n/3})$ $p$ is odd and $|\theta|=3$ 4.3
(7) $q^{1+2m}_{+}\cong p^{n+2mn}$ $p^{2mn^{2}}{:}({\mathrm{Sp}}(2m,q){:}{\Gamma\mathrm{L}}(1,q))$ $q=p^{n}$ and $p$ is odd 4.4
(8) $q^{1+2m}_{+}/U\cong p^{n_{0}+2mn}$ $p^{2mn_{0}n}{:}({\mathrm{Sp}}(2m,q){:}{\Gamma\mathrm{L}}(1,q)_{U})$ Example 5.3 5.4
In Table 1, groups $A_{p}(n,\theta)$’s are known as Suzuki $p$-group of type A (see Example 2.4); the group $P(\epsilon)$ was first given by Dornhoff [12] (see Theorem 2.6); and $q^{1+2m}_{+}$ is a generalization of $p^{1+2m}_{+}$, the extraspecial $p$-group of exponent $p$ with odd prime $p$. We now explain how to generalize $p^{1+2m}_{+}$ to $q^{1+2m}_{+}$.
Let $q=p^{n}$ for an odd prime $p$. For convenience, we denote a Sylow $p$-subgroup of a group $X$ by $X_{p}$. It is known that $p^{1+2}_{+}\cong{\mathrm{SL}}(3,p)_{p}$ and $p^{1+2m}_{+}$ is the central product of $m$-copies of $p^{1+2}_{+}$. A natural way to generalize such groups is defining $q^{1+2}_{+}$ to be ${\mathrm{SL}}(3,q)_{p}$ and defining $q^{1+2m}_{+}$ to be the central product of $m$-copies of $q^{1+2}_{+}$ (also see Definition 2.17). A classical result of Winter [31] says that ${\mathrm{Out}}(p^{1+2m}_{+})\cong\mathrm{CSp}(2m,p)\cong{\mathrm{Sp}}(2m,p){:}{\mathbb{Z}}_{p-1}$. In Lemma 4.4, we prove that ${\mathrm{Aut}}(q^{1+2m}_{+})\cong p^{2mn^{2}}{:}{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$, which shows some similarities between $p^{1+2m}_{+}$ and $q^{1+2m}_{+}$. Thus, we name the group $q^{1+2m}_{+}$ the extraspecial $q$-group.
For odd prime $p$, we can observe from Table 1 that a non-abelian $3$-orbit $p$-group is isomorphic to either $A_{p}(n,\theta)$ with $|\theta|=3$ or a quotient group of $q^{1+2m}_{+}$ for some $q=p^{n}$. The following corollary is a criterion for deciding the quotient group $q^{1+2m}_{+}/U$ to be $3$-orbit.
Corollary 1.1.
Let $N:=q^{1+2m}_{+}$, and let $H:={\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}\cong{\Gamma\mathrm{L}}(1,q)$. Then $N/U$ is a $3$-orbit group with $U<{\mathbf{Z}}(N)$ if and only if $H_{U}$ acts transitively on non-identity elements of ${\mathbf{Z}}(N)/U\cong{\mathbf{Z}}(N/U)$.
Additionally, we obtain some isomorphisms among some $p$-groups related to extraspecial $q$-groups, which is independently interesting. We remark that an ${\mathbb{F}}_{p}$-subspace $U$ of $V={\mathbb{F}}_{p^{n}}^{+}$ is called a subfield hyperplane with respect to subfield ${\mathbb{F}}_{p^{d}}$ for $d\mid n$ if and only if $U$ is a hyperplane of $V$ as ${\mathbb{F}}_{p^{d}}$-spaces (also see Definition 5.1).
Theorem C.
Let $q=p^{n}$ for odd prime $p$ and $n\geqslant 1$. Then the following statements hold.
  1. (i)
    Let $U<{\mathbf{Z}}(q^{1+2m}_{+})\cong{\mathbb{F}}_{q}^{+}$. Then $U$ is a subfield hyperplane of ${\mathbf{Z}}(q_{+}^{1+2m})\cong{\mathbb{F}}_{q}^{+}$ with respect to ${\mathbb{F}}_{q_{0}}$ if and only if
    $q^{1+2m}_{+}/U\cong(q_{0})^{1+2mn/d}_{+},$
    where $q_{0}=p^{d}$. In particular, for any subgroup $U<{\mathbf{Z}}(q^{1+2m}_{+})$ of order $p^{n-1}$, we have
    $q^{1+2m}_{+}/U\cong p^{1+2mn}_{+}.$
  2. (ii)
    $q^{1+2}_{+}\cong{\mathrm{SL}}(3,q)_{p}\cong{\mathrm{SU}}(3,q)_{p}$.
  3. (iii)
    $q^{1+2m}_{+}\cong\textbf{O}_{p}(G_{\alpha})$, where $G_{\alpha}$ is the stabilizer of ${\mathrm{PSp}}(2m+2,q)$ acting naturally on $1$-dimensional subspaces of ${\mathbb{F}}_{q}^{2m+2}$.
A transitive permutation group $G$ on a set ${\Omega}$ is said to be of rank $r$ if $G$ has exactly $r$ orbits on ${\Omega}\times{\Omega}$. The study of rank $3$ groups dates back to work of G. Higman [15] in 1964. Note that the holomorph ${\mathrm{Hol}}(N)=N{:}{\mathrm{Aut}}(N)$ of a group $N$ has rank $r$ if and only if $N$ is an $r$-group. Hence, our classification of $3$-orbit groups provides new families of rank $3$ groups.
Corollary 1.2.
The holomorph ${\mathrm{Hol}}(N)\cong N{:}{\mathrm{Aut}}(N)$ is a finite rank $3$ group if and only if $N$ is one of groups in Table 1.
The layout of the paper is as follows. In Section 2, we introduce some finite $3$-orbit groups, and then figure out the relations between representation theory and special $p$-groups. The proof of Theorem A will be given in Section 3. Section 4 contains constructions and proofs of properties of groups in rows 1-7 in Table 1. We will discuss the relations between subfield hyperplanes and groups in the 8-th row in Table 1 in Section 5, and then we complete proofs of Theorems B and C.
Acknowledgments
This work was supported by NNSFC grant no. 11931005. The authors thank to Eamonn A. O’Brien for his advice on Magma.
2. $3$-orbit groups and special $p$-groups
In this section, we draw some preliminary properties and results on $3$-orbit groups. As observed before, a finite $3$-orbit group is either a $p$-group of exponent at most $p^{2}$, or a $\{p,q\}$-group for two primes $p,q$.
2.1. Some known finite $3$-orbit groups
Finite $3$-orbit $\{p,q\}$-groups are classified in [10, 12, 20, 24], and are described in the following Example 2.1. Recall that a primitive prime divisor of $a^{k}-1$ is a prime $r$ such that $r\mid(a^{k}-1)$ but $r\nmid(a^{i}-1)$ for any $11\leqslant i\leqslant k-$ 1.
Example 2.1.
Let $p$ and $q$ be two distinct primes such that $q$ is a primitive prime divisor of $p^{q-1}-1$. Set $X={\mathrm{A}\Gamma\mathrm{L}}(n,p^{q-1})=P{:}H$ with $P\cong{\mathbb{Z}}_{p}^{n(q-1)}$ and $H\cong{\Gamma\mathrm{L}}(n,p^{q-1})$. Then $H$ has a normal cyclic subgroup $L\cong{\mathbb{Z}}_{p^{q-1}-1}$ which is the center of ${\mathrm{GL}}(n,p^{q-1})$. Noting that $q\mid p^{q-1}-1$, let $Q$ be the unique cyclic subgroup of $L$ of order $q$. Then the $\{p,q\}$-group
$N:=P{:}Q\cong{\mathbb{Z}}_{p}^{n(q-1)}{:}{\mathbb{Z}}_{q}$
is a normal subgroup of $X=P{:}H={\mathrm{A}\Gamma\mathrm{L}}(n,p^{q-1})$.
We make the following observations on the group $N:=P{:}Q$.
  1. (i)
    Viewing $P\cong{\mathbb{Z}}_{p}^{n(q-1)}$ as a vector space, $Q$ acting on $P$ is a scalar multiplication by definition, and so $P$ is a homogeneous ${\mathbb{F}}_{p}Q$-module. Thus, $N$ is a homogeneous Frobenius group, and ${\mathrm{Aut}}(N)\cong X={\mathrm{A}\Gamma\mathrm{L}}(n,p^{q-1})$, by [29, Theorem 1.1].
  2. (ii)
    Elements of $N$ of order $p$ are conjugate under ${\Gamma\mathrm{L}}(n,p^{q-1})$. Sylow $q$-subgroups of $N$ are conjugate. The field automorphism $\phi$ of ${\mathbb{F}}_{p^{q-1}}$ acts on $Q$ faithfully, and so non-identity elements of $Q$ are conjugate under $\langle\phi\rangle$. Thus, elements in $N$ of order $q$ are conjugate under ${\mathrm{Aut}}(N)$, and $N=P{:}Q$ is a $3$-orbit group.
By [10, 12, 20, 24], each $3$-orbit group of order divisible by primes $p,q$ is isomorphic to a homogeneous Frobenius group ${\mathbb{Z}}_{p}^{n(q-1)}{:}{\mathbb{Z}}_{q}$, as described in Example 2.1.
Lemma 2.2.
A finite group $3$-orbit group $N$ is not a $p$-group if and only if $N\cong{\mathbb{Z}}_{p}^{n(q-1)}{:}{\mathbb{Z}}_{q}$, as described in Example 2.1. In this case, ${\mathrm{Aut}}(N)\cong{\mathrm{A}\Gamma\mathrm{L}}(n,p^{q-1})$.
It is known and easily shown that each homocyclic group ${\mathbb{Z}}_{p^{2}}^{n}$ is a $3$-orbit group, and
${\mathrm{Aut}}({\mathbb{Z}}_{p^{2}}^{n})={\mathrm{GL}}(n,{\mathbb{Z}}/p^{2}{\mathbb{Z}})\cong p^{n^{2}}.{\mathrm{GL}}(n,p).$
Conversely, every finite abelian $3$-orbit group is homocyclic of exponent $p^{2}$.
Lemma 2.3.
A finite abelian group $N$ is a $3$-orbit group if and only if $N\cong{\mathbb{Z}}_{p^{2}}^{n}$ for some prime $p$ and integer $n$.
A non-abelian $2$-group with more than one involutions is called a Suzuki $2$-group if it has a cyclic group of automorphisms acting transitively on involutions. D. Higman classified all Suzuki $2$-group and divided them into four types A-D in [16]. The definition of each type of Suzuki $2$-group can be extended to odd primes, which are called Suzuki $p$-groups. The following is the definition of Suzuki $p$-group of type A, also see [2, Section 46].
Example 2.4.
Let $\theta\in{\mathrm{Gal}}({\mathbb{F}}_{p^{n}}/{\mathbb{F}}_{p})$. For $a,b\in{\mathbb{F}}_{p^{n}}$, define
$A_{p}(n,\theta):=\langle M(a,b,\theta)\mid a,b\in{\mathbb{F}}_{p^{n}}\rangle,$
where $M(a,b,\theta)$ is a matrix of the form
$M(a,b,\theta)=\begin{pmatrix}1&a&b\\ 0&1&a^{\theta}\\ 0&0&1\end{pmatrix}.$
We remark that groups $A_{2}(n,\theta)$ with $|\theta|\neq 1$ of odd order form the class of Suzuki $2$-groups of type A given by D. Higman. We collect some properties of $A_{p}(n,\theta)$ blow.
Proposition 2.5.
Let $N:=A_{p}(n,\theta)$. Then the following statements hold.
  1. (i)
    [22, Lemma 4.2] When $p=2$ and $|\theta|>1$ is odd, $N$ is a $3$-orbit $2$-group with ${\mathrm{Aut}}(N)\cong 2^{n^{2}}{:}{\Gamma\mathrm{L}}(1,2^{n})$.
  2. (ii)
    Let $\lambda$ be a generator of ${\mathbb{F}}_{p^{n}}$, and let $D=\mathrm{diag}(\lambda^{-1},1,\lambda^{\theta})$. Then $D$ induces an an automorphism $\xi$ of $N$ such that
    $\xi:M(a,b,\theta)\mapsto D^{-1}M(a,b,\theta)D=M(a\lambda,b\lambda\lambda^{\theta},\theta).$
We remark that $A_{2}(2f+1,\theta)$ for $x^{\theta}=x^{2^{f+1}}$ is isomorphic to a Sylow $2$-subgroup of Suzuki group ${\mathrm{Sz}}(2^{2f+1})$, see [8, Proposition 13.4(vi)]. Hence, Sylow $2$-subgroups of ${\mathrm{Sz}}(2^{2f+1})$ are $3$-orbit $2$-groups.
Finite $3$-orbit $2$-groups are classified independently by Bors and Glasby [4], and the authors [22]. We record the results blow.
Theorem 2.6.
Suppose that $N$ is a non-abelian $3$-orbit $2$-group. Then one of the following statement holds.
  1. (i)
    $N\cong A_{2}(n,\theta)$ for $|\theta|\neq 1$ of odd order, and ${\mathrm{Aut}}(N)\cong 2^{n^{2}}{:}{\Gamma\mathrm{L}}(1,2^{n})$.
  2. (ii)
    $N\cong{\mathrm{SU}}(3,2^{n})_{2}$ and ${\mathrm{Aut}}(N)\cong 2^{2n^{2}}{:}{\Gamma\mathrm{L}}(1,2^{2n})$.
  3. (iii)
    $N\cong P(\epsilon)$ and ${\mathrm{Aut}}(N)\cong 2^{18}{:}(7{:}9)$, where
    $P(\epsilon)=$ $\langle x_{i},z_{j}\mbox{ for $1\leqslant i\leqslant 6$ and $1\leqslant j\leqslant 3$ }|z_{j}^{2}=[x_{i},z_{j}]=[z_{k},z_{\ell}]=1\mbox{ foreach $i,j,k,\ell$,}$
    $x_{1}^{2}=x_{3}^{2}=z_{2},\ x_{2}^{2}=z_{2}z_{3},\ x_{4}^{2}=z_{3},\ x_{5}^{2}=z_{1}z_{2}z_{3},\ x_{6}^{2}=z_{3},$
    $[x_{1},x_{2}]=[x_{3},x_{5}]=[x_{3},x_{6}]=z_{1}z_{2},\ [x_{1},x_{3}]=z_{1}z_{3},\ [x_{1},x_{4}]=z_{3},$
    $[x_{1},x_{5}]=[x_{3},x_{4}]=[x_{5},x_{6}]=z_{2},\ [x_{1},x_{6}]=1,\ [x_{2},x_{6}]=z_{1}z_{2}z_{3},$
    $[x_{2},x_{3}]=[x_{2},x_{4}]=[x_{4},x_{6}]=z_{1},\ [x_{2},x_{5}]=[x_{4},x_{5}]=z_{2}z_{3}.\rangle$
2.2. Special $p$-groups of exponent $p$
Let $N$ be a non-abelian $3$-orbit $p$-group with $p$ an odd prime. Then non-identity elements in the center ${\mathbf{Z}}(N)$ form an automorphism orbit, and elements in $N\setminus{\mathbf{Z}}(N)$ form a different automorphism orbit. It follows that both ${\mathbf{Z}}(N)$ and $N/{\mathbf{Z}}(N)$ are elementary abelian groups, and ${\mathbf{Z}}(N)=\Phi(N)=N^{\prime}$. So $N$ is a special $p$-group of exponent $p$ or $p^{2}$. If the exponent of $N$ is $p^{2}$, then all elements of $N$ of order $p$ form an automorphism orbit, and so $N$ is abelian by [27, Corollary 3]. This leads to the following lemma.
Lemma 2.7.
Let $N$ be a non-abelian $3$-orbit $p$-group with $p$ an odd prime. Then $N$ is a special $p$-group with exponent $p$.
For a non-abelian special $p$-group $N$, ${\mathbf{Z}}(N)=N^{\prime}=\Phi(N)$ and $N/\Phi(N)$ are elementary abelian $p$-groups, and thus they can be viewed as vector spaces over field ${\mathbb{F}}_{p}$. Then the actions of $A:={\mathrm{Aut}}(N)$ on $\Phi(N)$ and $N/\Phi(N)$ are linear actions. It yields that $\Phi(N)$ and $N/\Phi(N)$ can be viewed as ${\mathbb{F}}_{p}A$-modules. For elements $x,y,z\in N$, we have $[x,yz]=[x,y]^{z}[x,z]=[x,y][x,z]$ as ${\mathbf{Z}}(N)=N^{\prime}$. Thus, the commutator operator [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] on $N$ resembles an alternating bilinear form. Let $V$ be an ${\mathrm{F}}G$-module. The exterior square $\Lambda_{\mathrm{F}}^{2}(V)$ is the quotient ${\mathrm{F}}G$-module $(V\otimes V)/S$, where $S$ is spanned by vectors $v\otimes v$ for all $v\in V$. The exterior square has a universal property associated with alternating bilinear form. The following lemma is given in [22, Lemma 3.1].
Lemma 2.8.
Let $N$ be a non-abelian special $p$-group with $p$ prime such that $N/{\mathbf{Z}}(N)\cong{\mathbb{Z}}_{p}^{n}$. Let $K$ and $L$ be the kernels of $A:={\mathrm{Aut}}(N)$ acting on $V:=N/{\mathbf{Z}}(N)$ and $M:={\mathbf{Z}}(N)$, respectively. Let $G=A/K$. Then the following statements hold.
  1. (i)
    $K$ is an elementary abelian $p$-group of order $|{\mathbf{Z}}(N)|^{n}$.
  2. (ii)
    $K\leqslant L$, and then both $V$ and $M$ are naturally ${\mathbb{F}}_{p}G$-modules.
  3. (iii)
    $N$ is an $r$-group with $r=o(V)+o(M)-1$, where $o(V)$ and $o(M)$ are numbers of $G$-orbits on $V$ and $M$, respectively.
  4. (iv)
    $M\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ for some ${\mathbb{F}}_{p}G$-submodule $W$ of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$.
Recall that, for an odd prime $p$, a non-abelian $3$-orbit $p$-group $N$ is a special $p$-group of exponent $p$. Let ${\mathcal{H}}_{n,p}$ denote the $n$-generator free group in the variety of groups that have exponent $p$ and nilpotency class $2$. Then $N$ is a quotient of ${\mathcal{H}}_{n,p}$ when $N/\Phi(N)\cong{\mathbb{Z}}_{p}^{n}$. We remark that ${\mathcal{H}}_{n,p}$ is the so-called the universal generalized Heisenberg group in [24], and it is a finite special $p$-group with $|{\mathbf{Z}}({\mathcal{H}}_{n,p})|=p^{\binom{n}{2}}$ (see [3, Page 107]). We formulate an equivalent definition below.
Definition 2.9.
Let $p$ be an odd prime, and let $n$ be a positive integer. Define
${\mathcal{H}}_{n,p}:=\langle\alpha_{1},...,\alpha_{n}|\alpha_{i}^{p}=[\alpha_{i},[\alpha_{j},\alpha_{k}]]=1\mbox{ for $i,j,k=1,...,n$}.\rangle$
Let ${\mathcal{H}}={\mathcal{H}}_{n,p}$ and let $W<{\mathbf{Z}}({\mathcal{H}})$. The automorphism group of ${\mathcal{H}}/W$ for $W<{\mathbf{Z}}({\mathcal{H}})$ is described in [13, Lemma 3] which is based on [26, Theorem 2.10]. We state the results below.
Lemma 2.10.
Suppose that $N={\mathcal{H}}/W$ where ${\mathcal{H}}:={\mathcal{H}}_{n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Identify $V:=N/{\mathbf{Z}}(N)$ with ${\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$, and identify ${\mathbf{Z}}(N)$ with ${\mathbf{Z}}({\mathcal{H}})/W$. Then the following statements hold.
  1. (i)
    $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{Aut}}(V)={\mathrm{GL}}(V)\cong{\mathrm{GL}}(n,p)$, and ${\mathbf{Z}}({\mathcal{H}})\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ as an ${\mathbb{F}}_{p}X$-module.
  2. (ii)
    ${\mathrm{Aut}}(N)^{V}\leqslant X$ is equal to $X_{W}$, the maximal subgroup of $X$ stabilizes the subspace $W$ of ${\mathbf{Z}}({\mathcal{H}})\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$.
Using the notations above, Lemma 2.8 (i) deduces that
${\mathrm{Aut}}({\mathcal{H}})=K.X\cong p^{\frac{n^{2}(n-1)}{2}}.{\mathrm{GL}}(V),$
where $K\cong p^{\frac{n^{2}(n-1)}{2}}$ is the kernel of $X$ acting on ${\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$. The following simple lemma shows that the group extension $K.X$ is split.
Lemma 2.11.
Suppose that $N={\mathcal{H}}/W$ where ${\mathcal{H}}:={\mathcal{H}}_{n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Let $V:=N/{\mathbf{Z}}(N)$, and let $K$ be the kernel of ${\mathrm{Aut}}(N)$ acting on $V$. Then the group extension ${\mathrm{Aut}}(N)=K.{\mathrm{Aut}}(N)^{V}$ is split.
Proof.
First, we consider $N={\mathcal{H}}$. Then $K\cong p^{\frac{n^{2}(n-1)}{2}}$ and ${\mathrm{Aut}}({\mathcal{H}})^{V}\cong{\mathrm{GL}}(V)\cong{\mathrm{GL}}(n,p)$ by Lemma 2.10. Recall that ${\mathcal{H}}=\langle\alpha_{1},...,\alpha_{n}\rangle$. Let $k\in{\mathbb{F}}_{p}^{\times}$ be a primitive element, then there exists $\psi\in{\mathrm{Aut}}({\mathcal{H}})$ of order $p-1$ such that
$\psi:\alpha_{i}\mapsto x_{i}\alpha_{i}^{k}\mbox{, for some $x_{i}\in{\mathbf{Z}}({\mathcal{H}})$ with $i=1,...,n$.}$
Since $K$ is the kernel of ${\mathrm{Aut}}({\mathcal{H}})$ acting on $V={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$, for each $\varphi\in K$ there exist $a_{1},...,a_{n}\in{\mathbf{Z}}({\mathcal{H}})$ such that
$\varphi:\alpha_{i}\mapsto a_{i}\alpha_{i},\mbox{ for }i=1,...,n.$
Note that $a_{i}$ is a product of some $[\alpha_{i},\alpha_{j}]$’s and
$[\alpha_{i},\alpha_{j}]^{\psi}=[\alpha_{i}^{\psi},\alpha_{j}^{\psi}]=[\alpha_{i},\alpha_{j}]^{k^{2}}.$
Thus, we have $a_{i}^{\psi}=a_{i}^{k^{2}}$ for each $i,...,n$. Let $s\in{\mathbb{F}}_{p}^{\times}$ such that $sk=1$. Then $\alpha_{i}^{\psi^{-1}}=x_{i}^{-s^{3}}\alpha_{i}^{s}$ and
$\alpha_{i}^{\psi^{-1}\varphi\psi}=(x_{i}^{-s^{3}}\alpha_{i}^{s})^{\varphi\psi}=(x_{i}^{-s^{3}}a_{i}^{s}\alpha_{i}^{s})^{\psi}=(a_{i}^{s})^{\psi}\alpha_{i}=a_{i}^{k}\alpha_{i}.$
It follows that $\varphi^{\psi}=\varphi^{k}$, and hence $C_{K}(\psi)=1$.
Let $\phi\in{\mathrm{Aut}}({\mathcal{H}})$. Then $[\phi K,\psi K]=1$ since $\langle\psi K\rangle={\mathbf{Z}}({\mathrm{Aut}}({\mathcal{H}})/K)\cong{\mathbb{Z}}({\mathrm{GL}}(V))$. Note that $k\neq 1$, there exists $t\in{\mathbb{F}}_{p}^{\times}$ such that $t(1-k)=1$. Let $\varphi=\left[\psi^{\phi^{-1}},\phi\right]^{t}\in K$, and let $\phi^{\prime}=\varphi\phi$. Then
$=[\varphi,\psi]^{\phi}[\phi,\psi]=\left((\varphi^{-1})^{\psi}\varphi\right)^{\phi}[\phi,\psi]=\left(\varphi^{1-k}[\phi,\psi]^{\phi^{-1}}\right)^{\phi}$
$=\left(\left[\psi^{\phi^{-1}},\phi\right]\left[\phi,\psi^{\phi^{-1}}\right]\right)^{\phi}=1.$
Thus, for any $\phi\in{\mathrm{Aut}}({\mathcal{H}})$, there exists $\phi^{\prime}\in{\mathrm{Aut}}({\mathcal{H}})$ such that $\phi^{\prime}K=\phi K$ and $[\phi^{\prime},\psi]=1$. It follows that
$|C_{{\mathrm{Aut}}({\mathcal{H}})}(\psi)|\geqslant|{\mathrm{Aut}}({\mathcal{H}})^{V}|=|{\mathrm{GL}}(V)|.$
Since $C_{K}(\psi)=1$ and ${\mathrm{Aut}}({\mathcal{H}})\cong K.{\mathrm{GL}}(V)$, we have $C_{{\mathrm{Aut}}({\mathcal{H}})}(\psi)\cong{\mathrm{GL}}(V)$ is a complement of $K$ in ${\mathrm{Aut}}({\mathcal{H}})$.
Now we consider the general case that $N={\mathcal{H}}/W$. Let $X$ be a complement of $K$ in ${\mathrm{Aut}}({\mathcal{H}})$. Then $X_{W}$ admits a group of automorphisms of $N$, and is a complement of $K$ in ${\mathrm{Aut}}(N)$ where $K$ is the kernel of ${\mathrm{Aut}}(N)$ acting on $V$. Therefore, the group extension ${\mathrm{Aut}}(N)=K.{\mathrm{Aut}}(N)^{V}$ is split as ${\mathrm{Aut}}(N)^{V}\cong X_{W}$ by Lemma 2.10 (ii). ∎
Every $g\in{\mathrm{GL}}(n,p)$ induces an automorphism $\sigma\in{\mathrm{Aut}}({\mathcal{H}})$ by Lemma 2.10 (i), where ${\mathcal{H}}={\mathcal{H}}_{n,p}$. Thus, we have a straightforward isomorphism ${\mathcal{H}}/W\cong{\mathcal{H}}/W^{\sigma}$ for any $W<{\mathbf{Z}}({\mathcal{H}})$, which yields the following lemma.
Lemma 2.12.
Let ${\mathcal{H}}:={\mathcal{H}}_{n,p}$, and let $W_{1},W_{2}$ be two subgroups of ${\mathbf{Z}}({\mathcal{H}})$. View $W_{1}$ and $W_{2}$ as two subspaces of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ where $V={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$. Then ${\mathcal{H}}/W_{1}\cong{\mathcal{H}}/W_{2}$ if there exists $g\in{\mathrm{GL}}(V)$ such that $W_{1}^{g}=W_{2}$ in $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$.
2.3. Extraspecial $q$-groups
Let $p$ be an odd prime. We introduce a new family of special $p$-groups for odd prime $p$ in this part. Recall that the extraspecial $p$-group of exponent $p$, denoted by $p^{1+2m}_{+}$, is the central product of $m$ copies of ${\mathrm{SL}}(3,p)_{p}$ for odd prime $p$. This leads us to start with the Sylow $p$-subgroups of ${\mathrm{SL}}(3,p^{n})$.
Definition 2.13.
Let $q=p^{n}$ for odd prime $p$. The Heisenberg group (over finite field ${\mathbb{F}}_{q}$), denoted by $q^{1+2}_{+}$, is a subgroup of ${\mathrm{SL}}(3,q)$ consisting of matrices of form $\begin{pmatrix}1&a&c\\ 0&1&b\\ 0&0&1\end{pmatrix}$ for all $a,b,c\in{\mathbb{F}}_{q}$.
We remark that the Heisenberg group is a non-abelian special $p$-group. When $p=q$, it is exactly the extraspecial $p$-group of plus type, and our notation fits $p^{1+2}_{+}$.
Lemma 2.14.
Let $N:=q^{1+2}_{+}$ for $q=p^{n}$ where $p$ is an odd prime. Then $p^{2n^{2}}{:}{\Gamma\mathrm{L}}(2,q)\lesssim{\mathrm{Aut}}(N)$ and $N$ is a $3$-orbit group.
Proof.
For $x,y,z\in{\mathbb{F}}_{q}$, let
$A(x):=\begin{pmatrix}1&x&0\\ 0&1&0\\ 0&0&1\end{pmatrix},B(y):=\begin{pmatrix}1&0&0\\ 0&1&y\\ 0&0&1\end{pmatrix}\mbox{ and }C(z):=\begin{pmatrix}1&0&z\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
Then
  1. (1)
    $[A(x),B(y)]=C(xy)$, $[A(x),A(y)]=[B(x),B(y)]=1$; and
  2. (2)
    $[A(x_{1}),B(y_{1})][A(x_{2}),B(y_{2})]=[A(x_{1}+x_{2}),B(y_{1}+y_{2})]$ for $x_{1},x_{2},y_{1},y_{2}\in{\mathbb{F}}_{q}$.
Since $N$ is a special $p$-group of exponent $p$, it is isomorphic to ${\mathcal{H}}/W$ with ${\mathcal{H}}={\mathcal{H}}_{2n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Let $\alpha(x)$ and $\beta(y)$ be preimages of $A(x)$ and $B(y)$ in ${\mathcal{H}}$ for $x,y\in{\mathbb{F}}_{q}$, respectively. Hence, we have
${\mathcal{H}}=\langle\alpha(x),\beta(y):x,y\in{\mathbb{F}}_{q}\rangle.$
By Lemma 2.10 (i), $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)\cong{\mathrm{GL}}(2n,p)$ where $V:={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})=N/{\mathbf{Z}}(N)$. Note that $V$ can be viewed as a $2$-dimensional ${\mathbb{F}}_{q}$-space $V_{q}$ with basis $\alpha(1){\mathbf{Z}}({\mathcal{H}})$ and $\beta(1){\mathbf{Z}}({\mathcal{H}})$. Let $G\cong{\Gamma\mathrm{L}}(2,p^{n})$ be the subgroup of $X$ acting naturally on $V_{q}$. Now we prove that $W^{g}=W$ for $g\in G$, and then $G\lesssim{\mathrm{Aut}}(N)^{V}$ by Lemma 2.10 (ii). Note that each element in ${\mathbf{Z}}({\mathcal{H}})={\mathcal{H}}^{\prime}$ is a product of elements of forms $[\alpha(x),\alpha(y)]$, $[\beta(x),\beta(y)]$ and $[\alpha(x),\beta(y)]$ for $x,y\in{\mathbb{F}}_{q}$.
  1. (1)
    Since $[A(x),A(y)]=1$, we have $[\alpha(x),\alpha(y)]\in W$. Assume that $(\alpha(1){\mathbf{Z}}({\mathcal{H}}))^{g}=\alpha(a)\beta(b){\mathbf{Z}}({\mathcal{H}})$ for some $a,b\in{\mathbb{F}}_{q}$. Then
    $[\alpha(x),\alpha(y)]^{g}=[\alpha(ax)\beta(bx),\alpha(ay)\beta(by)].$
    We obtain that $[\alpha(x),\alpha(y)]^{g}\in W$ since
    $=[A(ax),B(by)][B(bx),A(ay)]$
    $=C(ax\cdot by-bx\cdot ay)=1.$
  2. (2)
    Similarly, $[\beta(x),\beta(y)]\in W$ and $[\beta(x),\beta(y)]^{g}\in W$ for any $x,y\in{\mathbb{F}}_{q}$ and $g\in G$.
  3. (3)
    We assume that $w\in W$ with $w=\prod_{i=1}^{k}[\alpha(x_{i}),\beta(y_{i})]$ for some $x_{i},y_{i}\in{\mathbb{F}}_{q}$. Notice that $\sum_{i=1}^{k}x_{i}y_{i}=0$ since $\prod_{i=1}^{k}[A(x_{i}),B(y_{i})]=C\left(\sum_{i=1}^{n}x_{i}y_{i}\right)$. Let $a,b,c,d\in{\mathbb{F}}_{q}$ such that
    $(\alpha(1){\mathbf{Z}}({\mathcal{H}}))^{g}=\alpha(a)\beta(b){\mathbf{Z}}({\mathcal{H}})\mbox{ and }(\beta(1){\mathbf{Z}}({\mathcal{H}}))^{g}=\alpha(c)\beta(d){\mathbf{Z}}({\mathcal{H}}).$
    By calculations, we have $w^{g}$ equals to
    $\prod_{i=1}^{k}[\alpha(x_{i}),\beta(y_{i})]^{g}=\prod_{i=1}^{k}[\alpha(x_{i}a)\beta(x_{i}b),\alpha(y_{i}c)\beta(y_{i}d)]=\prod_{i=1}^{k}s_{i}\cdot\prod_{i=1}^{k}t_{i}.$
    where $s_{i}=[\alpha(x_{i}a),\alpha(y_{i}c)][\beta(x_{i}b),\beta(y_{i}d)]$, and $t_{i}=[\alpha(x_{i}a),\beta(y_{i}d)][\beta(x_{i}b),\alpha(y_{i}c)]$. By (1) and (2), $s_{i}$’s are in $W$. The image of $\prod_{i=1}^{k}t_{i}$ in $N$ is equal to
    $\prod_{i=1}^{k}[A(x_{i}a),B(y_{i}d)][B(x_{i}b),A(y_{i}c)]=C\left((ad-bc)\cdot\sum_{i=1}^{k}x_{i}y_{i}\right).$
    Since $\sum_{i=1}^{k}x_{i}y_{i}=0$, the image of $\prod_{i=1}^{k}t_{i}$ in $N$ is trivial. Hence, the image of $w^{g}$ in $N$ is trivial, that is, $w^{g}\in W$.
Hence, $W^{G}=W$ and $G\lesssim{\mathrm{Aut}}(N)^{V}$ by Lemma 2.10 (ii). Thus, ${\mathrm{Aut}}(N)\geqslant K.G\cong p^{2n^{2}}.{\Gamma\mathrm{L}}(2,q)$ where $K$ is the kernel of ${\mathrm{Aut}}(N)$ acting on $N/{\mathbf{Z}}(N)$ by Lemma 2.8 (i). In particular, the extension $K.G$ is split by Lemma 2.11, and then $p^{2n^{2}}{:}{\Gamma\mathrm{L}}(2,q)\lesssim{\mathrm{Aut}}(N)$.
Note that there exists an automorphism $\xi$ of $N$ admitted by conjugating of diagonal matrix $\mathrm{diag}(\lambda,1,1)$ where $\lambda$ is a generator of ${\mathbb{F}}_{q}$. By definition we observe that $\langle\xi\rangle$ acts regularly on non-identity elements of ${\mathbf{Z}}(N)$. In addition, $G\cong{\Gamma\mathrm{L}}(2,q)$ acts transitive on non-zero vectors of $V=N/{\mathbf{Z}}(N)\cong{\mathbb{F}}_{p}^{2n}$. Hence ${\mathrm{Aut}}(N)$ acts transitively on non-identity elements of both $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$. By Lemma 2.8 (iii), $N$ is a $3$-orbit group. ∎
Note that ${\Gamma\mathrm{L}}(2,q)$ contains the Singer cyclic which is a cyclic group of order $q^{2}-1$ acting regularly on non-identity elements of ${\mathbb{F}}_{q}^{2}$. This yields a solvable subgroup of ${\mathrm{Aut}}(N)$ which acts on $N$ with $3$ orbits. Immediately, we obtain the follow lemma, which is Theorem C (ii).
Lemma 2.15.
Let $q=p^{n}$ with odd prime $p$. Then
$q^{1+2}_{+}\cong{\mathrm{SL}}(3,q)_{p}\cong{\mathrm{SU}}(3,q)_{p}.$
Proof.
Let $N:=q^{1+2}_{+}$, and let $A\cong p^{2n^{2}}{:}{\Gamma\mathrm{L}}(2,q)$ be the subgroup of ${\mathrm{Aut}}(N)$ given in Lemma 2.14. Then there exists $\varphi\in A$ of order $q^{2}-1$ such that $\varphi$ induces a transitive action on non-identity elements of $N/{\mathbf{Z}}(N)$. Let $\xi\in{\mathrm{Aut}}(N)$ induced by conjugating the diagonal matrix $\mathrm{diag}(\mu,1,1)$ for some generator $\mu$ of ${\mathbb{F}}_{q}^{\times}$. We remark that $\xi\in A$ and $\xi$ acts transitively on non-identity elements of ${\mathbf{Z}}(N)$. Let $K\cong p^{2n^{2}}$ be the kernel of $A$ acting on $N/{\mathbf{Z}}(N)$. Then $\langle K,\varphi,\xi\rangle$ is a solvable subgroup of $A\leqslant{\mathrm{Aut}}(N)$, and it acts on elements of $N$ with exactly $3$ orbits by Lemma 2.8 (iii). Thus, ${\mathrm{Aut}}(N)$ has a solvable subgroup with exactly $3$ orbits on $N$.
Note that $H:=\langle M(a,b)\mid a,b\in{\mathbb{F}}_{q^{2}},b+b^{q}+a^{1+q}=0\rangle$ is a Sylow $p$-subgroup of ${\mathrm{SU}}(3,q)$ (see [11, page 249]), where
$M(a,b)=\begin{pmatrix}1&a&b\\ 0&1&a^{q}\\ 0&0&1\end{pmatrix}.$
Let $D=\mathrm{diag}(\lambda^{-q},\lambda^{1-q},\lambda)$, where $\langle\lambda\rangle={\mathbb{F}}_{q^{2}}^{\times}$. Then $D^{-1}M(a,b)D=M(q\lambda,b\lambda^{q+1})$ and
$b\lambda^{q+1}+(b\lambda^{q+1})^{q}+(a\lambda)(a\lambda)^{q}=\lambda^{1+q}(b+b^{q}+a^{1+q}).$
Hence, $D$ induces an automorphism $\xi^{\prime}\in{\mathrm{Aut}}(H)$ such that $\langle\xi^{\prime}\rangle$ acting transitively on non-identity elements of both $H/{\mathbf{Z}}(H)$ and ${\mathbf{Z}}(H)$. By Lemma 2.8 (i) and (iii), ${\mathrm{Aut}}(H)$ has a solvable subgroup with exactly $3$ orbits on $H$.
Finite groups which have solvable groups of automorphisms with exactly $3$ orbits have been classified in [12]. In particular, there exists a unique such group (up to isomorphism) of order $p^{3n}$ with center of order $p^{n}$. Note that both $q^{1+2}_{+}$ and ${\mathrm{SU}}(3,q)_{p}$ has order $p^{2n}$ with center of order $p^{n}$. Therefore, we have $q^{1+2}_{+}\cong{\mathrm{SL}}(3,q)_{p}\cong{\mathrm{SU}}(3,q)_{p}$. ∎
With Lemma 2.15, we can reformulate Dornhoff’s Theorem given in [12].
Theorem 2.16 (Dornhoff).
Let $N$ be a finite group. If there exists a solvable subgroup of ${\mathrm{Aut}}(N)$ acting on $N$ with exactly $3$ orbits, then one of the following statement holds.
  1. (i)
    $N$ is an abelian group isomorphic to ${\mathbb{Z}}_{p}^{n}$ or ${\mathbb{Z}}_{p^{2}}^{n}$ for prime $p$.
  2. (ii)
    $N$ is a $\{p,q\}$-group given in Example 2.1.
  3. (iii)
    $N$ is a non-abelian special $2$-group isomorphic to one of $A_{2}(n,\theta)$ (with $|\theta|\neq 1$ odd), ${\mathrm{SU}}(3,2^{n})_{2}$ and $P(\epsilon)$.
  4. (iv)
    $N\cong q^{1+2}_{+}\cong{\mathrm{SU}}(3,q)_{p}\cong{\mathrm{SL}}(3,q)_{p}$ for odd prime power $q=p^{n}$.
  5. (v)
    $N\cong 3^{1+4}_{+}$.
Now we extend the definition of extraspecial $p$-groups.
Definition 2.17.
For $q=p^{n}$ with odd prime $p$. The extraspecial $q$-group is defined by
$q^{1+2m}_{+}:=q^{1+2}_{+}\circ\cdots\circ q^{1+2}_{+}=(q^{1+2}_{+})^{m}/C,$
where $C$ consists of elements $(X_{1},...,X_{m})\in{\mathbf{Z}}(q^{1+2}_{+})^{m}$ such that $X_{1}\cdots X_{m}=1$.
It is straightforward to see that $q^{1+2m}_{+}$ is a non-abelian special $p$-group of exponent $p$ such that ${\mathbf{Z}}(q^{1+2m}_{+})\cong{\mathbb{Z}}_{p}^{n}$ and $q^{1+2m}_{+}/{\mathbf{Z}}(q^{1+2m}_{+})\cong{\mathbb{Z}}_{p}^{2mn}$. Let $\pi_{i}$ be the nature homomorphism from $q^{1+2}_{+}$ to the image of the $i$-th component in $q^{1+2m}_{+}$ for $i=1,...,m$. Remark that each matrix $Z\in{\mathbf{Z}}(q^{1+2}_{+})$ satisfies $\pi_{i}(Z)=\pi_{j}(Z)$ for $i,j=1,...,m$. Then we define the following notations for $i=1,...,m$ and $a,b,c\in{\mathbb{F}}_{q}$:
$E_{i}(a):=\pi_{i}\begin{pmatrix}1&a&0\\ 0&1&0\\ 0&0&1\end{pmatrix}\mbox{, }F_{i}(b):=\pi_{i}\begin{pmatrix}1&0&0\\ 0&1&b\\ 0&0&1\end{pmatrix}\mbox{, and }Z(c)=\pi_{i}\begin{pmatrix}1&0&c\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
The following equalities can be verified by simple calculations on matrices.
Proposition 2.18.
Using notations defined above.
  1. (i)
    $\pi_{i}\begin{pmatrix}1&a&c\\ 0&1&b\\ 0&0&1\end{pmatrix}=Z(c)F_{i}(b)E_{i}(a)$
  2. (ii)
    $E_{i}(a)E_{i}(b)=E_{i}(a+b),\ F_{i}(a)F_{i}(b)=F_{i}(a+b)\mbox{ and }\ Z(a)Z(b)=Z(a+b)$;
  3. (iii)
    $[E_{i}(a),E_{j}(b)]=[E_{i}(a),F_{j}(b)]=[F_{i}(a),F_{j}(b)]=1$ for $i\neq j$;
  4. (iv)
    $[E_{i}(a),F_{i}(b)]=Z(ab)$;
  5. (v)
    $E_{i}(a)F_{i}(b)=F_{i}(b)E_{i}(a)[E_{i}(a),F_{i}(b)]=Z(ab)F_{i}(b)E_{i}(a)$.
With above properties, we obtain the following lemma about ${\mathrm{Aut}}(q^{1+2m}_{+})$.
Lemma 2.19.
Let $N:=q^{1+2m}_{+}$ with $q=p^{n}$ for odd prime $p$. Then ${\mathrm{Aut}}(N)$ has a subgroup $A$ such that $A^{V}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$ and $A^{{\mathbf{Z}}(N)}\cong{\Gamma\mathrm{L}}(1,q)$, where $V=N/{\mathbf{Z}}(N)$.
Proof.
Since $N$ is a non-abelian special $p$-group of exponent $p$, $N={\mathcal{H}}/W$ with ${\mathcal{H}}:={\mathcal{H}}_{2mn,p}$ and $W\leqslant\Phi({\mathcal{H}})$. Identify $V:=N/{\mathbf{Z}}(N)$ with ${\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$, and let $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)$ as in Lemma 2.10. For convenience, let $\overline{\alpha}:=\alpha{\mathbf{Z}}({\mathcal{H}})\in V$ for $\alpha\in{\mathcal{H}}$. Let $e_{i}(a)$ and $f_{j}(b)$ be preimages of $E_{i}(a)$ and $F_{i}(b)$ in ${\mathcal{H}}$, respectively. Hence, $V$ can be viewed as an ${\mathbb{F}}_{q}$-symplectic space $V_{q}$ equipped with the symplectic form ${\mathcal{B}}$ and a basis
$\overline{e_{1}(1)},...,\ \overline{e_{m}(1)},\ \overline{f_{1}(1)},...,\ \overline{f_{m}(1)}$
such that $\overline{e_{i}(1)}$ and $\overline{f_{i}(1)}$ form a hyperbolic pair for each $i$. Let $G$ be the subgroup of $X$ isomorphic to ${\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$ acting naturally as a semilinear conformal symplectic group on $V_{q}$. We claim that $w^{g}\in W$ for $g\in G$ and $w\in W$, then $G\lesssim{\mathrm{Aut}}(N)^{V}$ by Lemma 2.10 (ii). Note that there exist some $\alpha_{i},\beta_{i}\in{\mathcal{H}}$ such that $w=\prod_{i=1}^{k}[\alpha_{i},\beta_{i}]$. We prove the claim step by step.
Step 1: we show that the image of $[\alpha,\beta]$ in $N$ is $Z\left({\mathcal{B}}\left(\overline{\alpha},\overline{\beta}\right)\right)$ for any $\alpha,\beta\in{\mathcal{H}}$. Assume that
$\overline{\alpha}=\prod_{i=1}^{m}\overline{e_{i}(a_{i})f_{i}(b_{i})}\mbox{ and}\overline{\beta}=\prod_{i=1}^{m}\overline{e_{i}(c_{i})f_{i}(d_{i})},$
where $a_{i},b_{i},c_{i},d_{i}\in{\mathbb{F}}_{q}$. Then the image of $[\alpha,\beta]$ in $N$ equals to
$\left[\prod_{i=1}^{m}(E_{i}(a_{i})F_{i}(b_{i})),\prod_{i=1}^{m}(E_{i}(c_{i})F_{i}(d_{i}))\right]$ $=\prod_{i=1}^{m}[E_{i}(a_{i}),F_{i}(d_{i})][F_{i}(b_{i}),E_{i}(c_{i})]$
$=\prod_{i=1}^{m}Z(a_{i}d_{i}-b_{i}c_{i})=Z\left(\sum_{i=1}^{m}(a_{i}d_{i}-b_{i}c_{i})\right)$
$=Z\left({\mathcal{B}}\left(\overline{\alpha},\overline{\beta}\right)\right).$
Step 2: we claim that $w\in W$ if and only if $\sum_{i=1}^{k}{\mathcal{B}}\left(\overline{\alpha_{i}},\overline{\beta_{i}}\right)=0$. By the previous step, we obtain that the image of $w$ in $N$ is equal to $Z\left(\sum_{i=1}^{k}{\mathcal{B}}\left(\overline{\alpha_{i}},\overline{\beta_{i}}\right)\right)$. This yields the claim since $w\in W$ if and only if $w$ has a trivial image in $N$.
Step 3: we are now ready to show that $w^{g}\in W$. Note that $w^{g}=\prod_{i=1}^{k}[\alpha_{i}^{\widetilde{g}},\beta_{i}^{\widetilde{g}}]$ where $\widetilde{g}$ is a preimage of $g$ in ${\mathrm{Aut}}({\mathcal{H}})$. We remark that $\overline{\alpha^{\widetilde{g}}}=\overline{\alpha_{i}}^{g}$ for $\alpha\in{\mathcal{H}}$. We may assume that ${\mathcal{B}}(v^{g},w^{g})=\lambda{\mathcal{B}}(v,w)^{\sigma}$ for $v,w\in V$ where some $\lambda\in{\mathbb{F}}_{q}$ and $\sigma\in{\mathrm{Aut}}({\mathbb{F}}_{q})$. Hence, the image of $w^{g}=\prod_{i=1}^{k}[\alpha_{i}^{\widetilde{g}},\beta_{i}^{\widetilde{g}}]$ in $N$ is equal to
$\prod_{i=1}^{k}Z\left({\mathcal{B}}\left(\overline{\alpha_{i}^{\widetilde{g}}},\overline{\beta_{i}^{\widetilde{g}}}\right)\right)=Z\left(\sum_{i=1}^{k}{\mathcal{B}}\left(\overline{\alpha_{i}}^{g},\overline{\beta_{i}}^{g}\right)\right)=Z\left(\lambda\left(\sum_{i=1}^{k}{\mathcal{B}}(\overline{\alpha_{i}},\overline{\beta_{i}})\right)^{\sigma}\right)=Z(0).$
Thus, $w^{g}\in W$ for any $w\in W$ and $g\in G$. Hence, $G$ is a subgroup of ${\mathrm{Aut}}(N)^{V}$. By Lemmas 2.8 (i) and 2.11, the preimage $A$ of $G$ in ${\mathrm{Aut}}(N)$ is isomorphic to $p^{2mn^{2}}{:}{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$, and $A^{V}=G\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$.
It is clear that $V$ is the natural ${\mathbb{F}}_{p}G$-module. We now prove that ${\mathbf{Z}}(N)$ is the natural ${\mathbb{F}}_{p}{\Gamma\mathrm{L}}(1,q)$-module. For $g\in G$, we assume that ${\mathcal{B}}(v^{g},w^{g})=\lambda{\mathcal{B}}(v,w)^{\sigma}$ for any $v,w\in V_{q}$ with $\lambda\in{\mathbb{F}}_{q}$ and $\sigma\in{\mathrm{Aut}}({\mathbb{F}}_{q})$. Let $g_{N}$ be the automorphism of $N$ induced by $g$. Then $[e_{1}(a)^{g},f_{1}(1)^{g}]$ is a preimage of $Z(a)^{g_{N}}=[E_{1}(a),F_{1}(1)]^{g_{N}}$ in ${\mathcal{H}}$. Hence, we have
$Z(a)^{g}=Z\left({\mathcal{B}}\left(\overline{e_{1}(a)}^{g},\overline{f_{1}(1)}^{g}\right)\right)=Z(\lambda a^{\sigma}).$
This yields that the image of $G$ acting on ${\mathbf{Z}}(N)$ is isomorphic to ${\Gamma\mathrm{L}}(1,q)$. Thus, we have $A^{{\mathbf{Z}}(N)}=G^{{\mathbf{Z}}(N)}\cong{\Gamma\mathrm{L}}(1,q)$. ∎
3. Exterior Square of Transitive Linear Groups
Let $p$ be an odd prime, $V$ be a vector space over ${\mathbb{F}}_{p}$. To classify special $3$-orbit $p$-group, Lemma 2.8 (iii) and (iv) lead us to focus on the pairs $(G,W)$ such that $W<\Lambda_{{\mathbb{F}}_{p}}^{2}(V)$ is invariant under $G\leqslant{\mathrm{GL}}(V)$ and $G$ acts transitive on non-zero vectors of both $V$ and $\Lambda_{{\mathbb{F}}_{p}}^{2}(V)/W$.
We remark that a linear group $G\leqslant{\mathrm{GL}}(V)$ acting transitively on non-zero vectors is called a transitive linear group. The classification of transitive linear groups is a part of well-known classification of $2$-transitive groups. The following theorem can be concluded from [6, Table 7.3].
Theorem 3.1.
Let $G\leqslant{\mathrm{GL}}(n,p)$ be a transitive linear group with odd prime $p$. Then one of the following statements hold.
  1. (i)
    $G$ is solvable.
  2. (ii)
    $G^{(\infty)}$ is isomorphic to ${\mathrm{SL}}(m,p^{f})$ for $m\geqslant 2$ and $(m,p^{f})\neq(2,3)$; or isomorphic to ${\mathrm{Sp}}(2m,p^{f})$ for $m\geqslant 2$.
  3. (iii)
    $G^{(\infty)}\cong{\mathrm{SL}}(2,5)$ with $(n,p)\in\{(2,11),(2,19),(2,29),(2,59)\}$.
  4. (iv)
    $G^{(\infty)}$ is isomorphic to $2^{1+4}.{\mathrm{A}}_{5}$ for $(n,p)=(4,3)$; or isomorphic to ${\mathrm{SL}}(2,13)$ for $(n,p)=(6,3)$.
Let $U$ be an ${\mathrm{E}}G$-module corresponding to the homomorphism $\rho:G\rightarrow{\mathrm{GL}}(V)$. For $\phi\in{\mathrm{Gal}}({\mathrm{E}}/{\mathrm{F}})$, denote by $U^{\phi}$ the ${\mathrm{E}}G$-module such that $u^{g}:=(u^{\rho(g)})^{\phi}$ for $u\in U$ and $g\in G$. For an ${\mathrm{F}}G$-module $V$, $V^{\mathrm{E}}:=V\otimes_{\mathrm{F}}{\mathrm{E}}$ is naturally an ${\mathrm{E}}G$-module. The following lemma on decompositions of exterior squares is [22, Lemma 2.2].
Lemma 3.2.
Suppose that $V$ is an ${\mathrm{F}}G$-module and ${\mathrm{E}}$ is an extension field of ${\mathrm{F}}$. If $U$ is an ${\mathrm{E}}G$-module such that $U\cong V$ as ${\mathrm{F}}G$-modules, then the following statements hold.
  1. (i)
    $V^{\mathrm{E}}\cong\bigoplus_{i=1}^{f}U^{\phi^{i}}$ where $\phi$ is a generator of ${\mathrm{Gal}}({\mathrm{E}}/{\mathrm{F}})$ and $f=[{\mathrm{E}}:{\mathrm{F}}]$;
  2. (ii)
    $\Lambda^{2}_{\mathrm{F}}(V)=A\oplus\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}$ where
    1. (1)
      $A\cong\Lambda_{\mathrm{E}}^{2}(U)$ as ${\mathrm{F}}G$-modules;
    2. (2)
      $B_{j}\cong U\otimes_{{\mathrm{E}}}U^{\phi^{j}}$ as ${\mathrm{F}}G$-modules, when $21\leqslant j<f/$ 2;
    3. (3)
      $2\cdot B_{f/2}\cong U\otimes_{{\mathrm{E}}}U^{\phi^{f/2}}$ as ${\mathrm{F}}G$-modules, when $f$ is even.
The following lemma is deduced by [22, Lemma 2.6] based on Steinberg’s Twisted Tensor Product Theorem [28] (also see [18, Theorem 5.4.1]).
Lemma 3.3.
Suppose that ${\mathrm{SL}}(m,q)\leqslant G\leqslant{\mathrm{GL}}(m,q)$ or ${\mathrm{Sp}}(m,q)\leqslant G\leqslant{\mathrm{CSp}}(2m,q)$ with $q=p^{f}$. Let $V$ and $U$ be the natural ${\mathbb{F}}_{p}G$-module and the natural ${\mathbb{F}}_{q}G$-module, respectively. Then $U\cong V$ as ${\mathbb{F}}_{p}G$-modules, and $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}$ where
  1. (i)
    $A\cong\Lambda_{{\mathbb{F}}_{q}}^{2}(U)$ as ${\mathbb{F}}_{p}G$-modules;
  2. (ii)
    for $j<f/2$, $B_{j}\cong U\otimes_{{\mathbb{F}}_{q}}U^{\phi^{j}}$ as irreducible ${\mathbb{F}}_{p}G$-module, and $B_{j}$’s are pairwise non-isomorphic;
  3. (iii)
    when $f$ is even, $2\cdot B_{f/2}\cong U\otimes_{{\mathbb{F}}_{q}}U^{\phi^{f/2}}$ as ${\mathbb{F}}_{p}G$-modules, and $B_{f/2}$ is an irreducible ${\mathbb{F}}_{p}G$-module.
The exterior squares of natural modules for classical groups are well-known. We state the following useful results for ${\mathrm{SL}}(m,q)$ below.
Lemma 3.4 ([23, Theorem 2.2]).
Let $U={\mathbb{F}}_{q}^{m}$ with $q=p^{f}$, and let $G={\mathrm{SL}}(m,q)$. Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is an irreducible ${\mathbb{F}}_{q}G$-module, and it is also an irreducible ${\mathbb{F}}_{p}G$-module.
In particular, $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is isomorphic to the dual of $U$ when $m=3$.
With the decomposition given in Lemma 3.3, we obtain the following lemma.
Lemma 3.5.
Suppose that $G\leqslant{\mathrm{GL}}(n,p)$ has a normal subgroup $G_{0}\cong{\mathrm{SL}}(m,p^{f})$ with $p$ a prime, $n=mf$ and $m\geqslant 3$. Let $V$ be the natural ${\mathbb{F}}_{p}G$-module. Then $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus B$ for ${\mathbb{F}}_{p}G$-submodules $A$ and $B$ such that
  1. (i)
    $A\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(U)$ as ${\mathbb{F}}_{p}G_{0}$-modules where $U$ is the natural ${\mathbb{F}}_{q}G_{0}$-module; and
  2. (ii)
    $G$ is transitive on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ if and only if $W=B$ and $m=3$.
Proof.
By Lemma 3.3, we can decompose $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ to be a direct sum of ${\mathbb{F}}_{p}G_{0}$-submodules:
$\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}.$
Let $B=\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}$. Then $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus B$ and part (i) holds by Lemma 3.3. Now we prove that both $A$ and $B$ are ${\mathbb{F}}_{p}G$-submodules. For $g\in G$, $\langle B^{g},B\rangle/B$ is an ${\mathbb{F}}_{p}G_{0}$-submodule of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/B\cong A$. Since $A$ is an irreducible ${\mathbb{F}}_{p}G_{0}$-module by Lemma 3.4, either $B^{g}=B$ or $A$ is isomorphic to one of $B_{j}$’s for $1\leqslant j\leqslant\lfloor\frac{f}{2}\rfloor$ as each $B_{j}$ is an irreducible ${\mathbb{F}}_{p}G_{0}$-module by Lemma 3.3. We can obtain that $\dim A=\frac{m(m-1)f}{2}$ and $\dim B_{j}=\left\{\begin{aligned} &m^{2}f,j\neq f/2;\\ &m^{2}f/2,j=f/2.\end{aligned}\right.$. Hence, $\dim A\neq\dim B_{j}$ for each $j$, and thus $B^{g}=B$ for any $g\in G$. This yields that $B$ is an ${\mathbb{F}}_{p}G$-submodule. On the other hand, if $A^{g}\neq A$, then $A$ is isomorphic to one of $B_{j}$’s, which is impossible. Thus, $A$ is also an ${\mathbb{F}}_{p}G$-submodule.
If $W=B$ and $m=3$, then $A$ is isomorphic to the dual of $V$ as ${\mathbb{F}}_{p}G_{0}$-modules by Lemma 3.4. Hence, $G_{0}$ acts transitively on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/B\cong A$, and so does $G$. Conversely, suppose that $G$ acts transitively on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$. Let $K$ be the kernel of $G$ acting on $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$. Note that $A$ and $B_{j}$’s are non-trivial irreducible ${\mathbb{F}}_{p}G_{0}$-submodules. Thus, $K\cap{\mathbf{Z}}(G_{0})\leqslant{\mathbf{Z}}(G_{0})$ and $G/K$ is non-solvable. By Theorem 3.1, we obtain that $K\cap{\mathbf{Z}}(G_{0})=1$, $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W=\dim V$ and $G_{0}$ is transitive on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$. If $B\not\subseteq W$, then $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W\cong B_{j}$ as ${\mathbb{F}}_{p}G_{0}$-modules for some $21\leqslant j\leqslant f/$ 2. Since $m\geqslant 3$, we obtain that $\dim B_{j}\geqslant mf=\dim V$. Thus, $B\subseteq W$, and then $B=W$ and $A\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$. Since $\dim A=\dim V$, we have $m=3$ as desired. ∎
Results of exterior square of the natural modules of symplectic groups can be found in [7, 19, 25], and we state them below.
Lemma 3.6.
Let $X={\mathrm{CSp}}(2m,q)$ and $G={\mathrm{Sp}}(2m,q)\lhd X$ acing naturally on symplectic space $U={\mathbb{F}}_{q}^{2m}$ with $q=p^{f}$ for odd prime $p$. Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ has a maximal ${\mathbb{F}}_{q}G$-submodule $T$ such that
  1. (i)
    $\dim_{{\mathbb{F}}_{q}}T=\dim_{{\mathbb{F}}_{q}}\Lambda^{2}_{{\mathbb{F}}_{q}}(U)-1$;
  2. (ii)
    $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)/T$ is a faithful ${\mathbb{F}}_{p}(X/G)$-module.
Moerover, $T$ is indecomposable and one the following statements holds.
  1. (i)
    When $p\nmid m$, $T$ is an irreducible ${\mathbb{F}}_{p}G$-module.
  2. (ii)
    When $p\mid m$, $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is an indecomposable ${\mathbb{F}}_{q}G$-module, $T$ has a unique maximal ${\mathbb{F}}_{q}G$-submodule $T_{0}$ which is trivial, and $T/T_{0}$ is an irreducible ${\mathbb{F}}_{p}G$-module.
Hence, we can deal with $\Lambda_{{\mathbb{F}}_{p}}^{2}(V)$ for the symplectic case.
Lemma 3.7.
Suppose that $G\leqslant{\mathrm{GL}}(n,p)$ has a normal subgroup $G_{0}\cong{\mathrm{Sp}}(2m,q)$ with $p$ an odd prime, $q=p^{f}$, $n=2mf$ and $(2m,q)\neq(2,3)$. Let $V$ be the natural ${\mathbb{F}}_{p}G$-module. Then $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus B$ for ${\mathbb{F}}_{p}G$-submodules $A$ and $B$ and the following statements hold.
  1. (i)
    $A\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(U)$ as ${\mathbb{F}}_{p}G_{0}$-modules where $U$ is the natural ${\mathbb{F}}_{q}G_{0}$-module.
  2. (ii)
    $A$ has an ${\mathbb{F}}_{p}G$-submodule $T$ such that $\dim A/T=f$ and $G_{0}$ is the kernel of $G$ acting on $A/T$.
  3. (iii)
    $S:=T\oplus B$ is an ${\mathbb{F}}_{p}G$-submodule and $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W\leqslant f$ if and only if $S\subseteq W$.
  4. (iv)
    $G$ is transitive on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ if and only if $S\subseteq W$ and $G/G_{0}\lesssim{\Gamma\mathrm{L}}(1,q)$ acting transitively on $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$.
Proof.
By Lemma 3.3, we can decompose $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ to be a direct sum of ${\mathbb{F}}_{p}G_{0}$-submodules:
$\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}.$
Let $B=\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}$. Then $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus B$ and part (i) holds by Lemma 3.3. Now we show that $A$ and $B$ are ${\mathbb{F}}_{p}G$-submodules. Note that if $B^{g}\neq B$ for some $g\in G$, then $A$ has an irreducible ${\mathbb{F}}_{p}G_{0}$-submodule whose dimension equals to $\dim B_{j}$ for some $1\leqslant j\leqslant\lfloor\frac{f}{2}\rfloor$. By Lemma 3.6, irreducible ${\mathbb{F}}_{p}G_{0}$-submodules of $A$ has dimension equal to one of $\left(\binom{2m}{2}-1\right)f$ and $\left(\binom{2m}{2}-2\right)f$, and none of them equal to $\dim B_{j}$ for any $j$. Hence, $B^{g}=B$ for any $g\in G$ and $B$ is an ${\mathbb{F}}_{p}G$-module. Similarly, we obtain that $A$ is also an ${\mathbb{F}}_{p}G$-submodule. By Lemma 3.6, $A$ has a maximal ${\mathbb{F}}_{q}G_{0}$-submodule $T$ such that $\dim_{{\mathbb{F}}_{p}}A/T=f$. If $T^{g}\neq T$, then $\langle T,T^{g}\rangle/T$ is a non-trivial $F_{p}G_{0}$=submodule of $A/T$, which is a contradiction. Thus, $T$ is an ${\mathbb{F}}_{p}G$-submodule of $A$. Let $K$ be the kernel of $G$ acting on $A/T$. Then $G_{0}=K\cap{\mathrm{CSp}}(2m,q)$ by Lemma 3.6. Since $G/G_{0}\leqslant{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)/{\mathrm{Sp}}(2m,q)\cong{\Gamma\mathrm{L}}(1,q)$, we have $K=G_{0}$. Hence, $G_{0}$ is the kernel of $G$ actign on $A/T$, as in part (ii).
Clearly that $S=T\oplus B$ is an ${\mathbb{F}}_{p}G$-submodule and $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/S=\dim A/T=f$. Hence $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W\leqslant f$ if $S\subseteq W$. We assume that $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W\leqslant f$. Note that each $B_{j}$ is an irreducible ${\mathbb{F}}_{p}G_{0}$-submodule with dimension larger than $f$, which yields that $B\subseteq W$. If $T$ is not an ${\mathbb{F}}_{p}G$-submodule of $W$, then $T\cap W$ is an ${\mathbb{F}}_{p}G$-submodule of codimension no more than $f$. When $p\nmid m$, $T$ is an irreducible ${\mathbb{F}}_{p}G_{0}$-module by Lemma 3.6, which yields a contradiction. When $p\mid m$, $T\cap W\subseteq T_{0}$ given in Lemma 3.6 where $\dim T_{0}=f$, which is also a contradiction to $\dim T/(T\cap W)\leqslant f$. Thus, we obtain that $T\subseteq W$ and $S:=T\oplus B\subseteq W$, as in part (iii).
Assume that $G$ is transitive on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$. If $G_{0}$ is not in the kernel of $G$ acting on $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$, then $G$ induces a non-solvable permutation group acting on $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$. In this case, $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ is an irreducible ${\mathbb{F}}_{p}G_{0}$-module with dimension equal to one of $n,2,4$ and $6$ by Theorem 3.1. We obtain that $m=1$, $2\mid f$ and $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W\cong B_{f/2}$ as ${\mathbb{F}}_{p}G_{0}$-modules. In this case, the induced permutation group of $G_{0}$ on $B_{f/2}$ is isomorphic to $\Omega^{-}_{4}(p^{f/2})$ (see [18, page 45]), which is not transitive on non-zero vectors of $B_{f/2}$. Thus, $G_{0}$ acts trivially on $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ and $B\subseteq W$. If $T$ is not an ${\mathbb{F}}_{p}G$-submodule of $W$, then $T/(T\cap W)$ is a trivial ${\mathbb{F}}_{p}G_{0}$-module, which is impossible by Lemma 3.6. We deduce that $S=B\oplus T\subseteq W$ and $G/G_{0}$ acting transitively on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ since $G_{0}$ acting trivially on $A/T$. This also yields the converse direction, and part (iv) holds. ∎
Now we are ready to prove Theorem 3.1.
Proof of Theorem 3.1.
Note that $G$ is isomorphic one of groups in Theorem 3.1. Using Magma [5], we obtain that if $G$ is one of the groups listed in Theorem 3.1 (iii) and (iv), then $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W=1$, as in case (i). Thus, we focus on groups $G$ with $G_{0}:=G^{(\infty)}$ isomorphic to one of ${\mathrm{SL}}(m,p^{f})$ ($m\geqslant 3$ and $n=mf$) and ${\mathrm{Sp}}(2m,p^{f})$ ($(m,p^{f})\neq(1,3)$ and $n=2mf$).
Assume that $G_{0}\cong{\mathrm{SL}}(m,p^{f})$ with $m\geqslant 3$ and $n=mf$. Then case (ii) holds by Lemma 3.5 (ii). Assume that $G_{0}\cong{\mathrm{Sp}}(2m,p^{f})$ with $(m,p^{f})\neq(1,3)$ and $n=2mf$. Lemma 3.7 (iv) deduces case (iii), and we complete the proof. ∎
4. $3$-orbit $p$-groups for odd prime $p$
Let $N$ be a finite non-abelian $3$-orbit $p$-group with $p$ an odd prime. Assume that $N/\Phi(N)\cong{\mathbb{Z}}_{p}^{n}$. Then $N$ is a quotient group ${\mathcal{H}}/W$ where ${\mathcal{H}}:={\mathcal{H}}_{n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Identify $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$ with ${\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$ and ${\mathbf{Z}}({\mathcal{H}})/W$, respectively. Let $V:=N/{\mathbf{Z}}(N)$ and $M:={\mathbf{Z}}(N)$. Recall Lemmas 2.8 (iii) and 2.10 that
  1. (i)
    $W$ is an ${\mathbb{F}}_{p}G$-submodule of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$;
  2. (ii)
    $G:={\mathrm{Aut}}(N)^{V}$ is the maximal subgroup of $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)$ in the sense that $W^{G}=W$; and
  3. (iii)
    $G$ acts transitively on non-zero vectors of both $V$ and $M\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$.
We may assume that $G$ is non-solvable due to Theorem 2.16. If $\dim M=1$, then $N$ is an extraspecial $p$-group, and hence $N\cong p^{1+2m}_{+}$ for some $m\geqslant 1$. Thus, we mainly focus on cases that $G^{(\infty)}$ is isomorphic to ${\mathrm{SL}}(3,p^{f})$ or ${\mathrm{Sp}}(2m,p^{f})$ by Theorem A.
4.1. ${\mathrm{SL}}(3,q)\lhd G$
Assume that $G_{0}=G^{(\infty)}\cong{\mathrm{SL}}(3,q)$ with $q=p^{f}$. We will show that such $3$-orbit group is uniquely determined by its order, and it is isomorphic to $A_{p}(n,\theta)$ with $|\theta|=3$ given in Example 2.4, as in the 6-th row of Table B.
First, we show the existence and the uniqueness of non-abelian $3$-orbit $p$-group $N$ with $|N|=|{\mathbf{Z}}(N)|^{2}$.
Lemma 4.1.
Let $p$ be an odd prime. Then the following statements hold.
  1. (i)
    If $N$ is a $3$-orbit $p$-group such that $|N|=|{\mathbf{Z}}(N)|^{2}$ then $|{\mathbf{Z}}(N)|=q^{3}$ for some $q=p^{f}$ and ${\mathrm{SL}}(3,q)\lhd{\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$.
  2. (ii)
    When $|N|=q^{6}$ for some $q=p^{f}$, there exists a unique $3$-orbit $p$-group such that $|{\mathbf{Z}}(N)|=q^{3}$.
  3. (iii)
    Let $N$ be the $3$-orbit $p$-group with $|N|=|{\mathbf{Z}}(N)|^{2}=q^{6}$ and $q=p^{f}$, then ${\mathrm{Aut}}(N)\cong p^{9f^{2}}{:}{\Gamma\mathrm{L}}(3,q)$.
Proof.
Assume that $N$ is a $3$-orbit $p$-group such that $|N|=|{\mathbf{Z}}(N)|^{2}$. By Theorem 2.16, ${\mathrm{Aut}}(N)$ must be non-solvable. Let $V:=N/{\mathbf{Z}}(N)$, $M:={\mathbf{Z}}(N)$ and $G:={\mathrm{Aut}}(N)^{V}$. Then $G$ acts transitively on non-zero vectors of both $V$ and $M$ by Lemma 2.8 (iii); and $M$ is isomorphic to $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ as ${\mathbb{F}}_{p}G$-modules by Lemma 2.8 (iv). Since $|M|=|V|$, we obtain that $G^{(\infty)}\cong{\mathrm{SL}}(3,q)$ and $|M|=|V|=q^{3}$ by Theorem A, as in part (i). Let ${\mathcal{H}}:={\mathcal{H}}_{n,p}$ with $n=3f$, and let $V:={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$. Then $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)\cong{\mathrm{GL}}(n,p)$ and ${\mathbf{Z}}({\mathcal{H}})\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ as ${\mathbb{F}}_{p}X$-modules by Lemma 2.10 (i). Let $G\leqslant X$ such that $G\cong{\Gamma\mathrm{L}}(3,q)$. Then ${\mathbf{Z}}({\mathcal{H}})\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ can be decomposed into $A\oplus B$ as ${\mathbb{F}}_{p}G$-modules such that $G$ acts transitively on non-zero vectors of ${\mathbf{Z}}({\mathcal{H}})/B\cong A$ by Lemma 3.5 (ii). Let $N:={\mathcal{H}}/B$. Then $G\leqslant{\mathrm{Aut}}(N)^{V}$ by Lemma 2.10 (ii), and $G$ acts transitively on non-identity elements of both $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$. Hence, $N$ is a $3$-orbit group by Lemma 2.8 (iii). This proves the existence in part (ii).
Now we prove the uniqueness in part (ii) by assuming that $N_{1}={\mathcal{H}}/W_{1}$ is another desired $3$-orbit $p$-group. Then part (i) deduces that ${\mathrm{SL}}(3,q)\lhd G_{1}:={\mathrm{Aut}}(N_{1})^{V}$. Note that there exists $x\in X$ such that $G_{1}^{x}\leqslant G$, and then $N_{1}\cong{\mathcal{H}}/W^{x}$ by Lemma 2.12. Thus, we may assume that $G_{1}\leqslant G$. Let $G_{0}=G^{(\infty)}=G_{1}^{(\infty)}\cong{\mathrm{SL}}(3,q)$. Then $W_{1}$ and $B$ are ${\mathbb{F}}_{p}G_{0}$-submodules of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ such that $G_{0}$ acts transitively on non-zero vectors of the corresponding quotient modules. By Lemma 3.5, we obtain that $W_{1}=B$, and therefore $N_{1}=N$.
Let $N={\mathcal{H}}/B$ as defined above. Then ${\mathrm{SL}}(3,q)\lhd{\mathrm{Aut}}(N)^{V}$ by part (i), and hence $G={\mathrm{Aut}}(N)^{V}\leqslant{\Gamma\mathrm{L}}(3,q)$. Note that $B^{g}=B$ for any $g\in{\Gamma\mathrm{L}}(3,q)$ by Lemma 3.5 (ii). Thus, we obtain that $G={\Gamma\mathrm{L}}(3,q)$ and ${\mathrm{Aut}}(N)\cong p^{n^{2}}{:}{\Gamma\mathrm{L}}(3,q)$ by Lemma 2.8 (i) and 2.11, as in part (iii). ∎
Recall that $|A_{p}(n,\theta)|=|{\mathbf{Z}}(A_{p}(n,\theta))|^{2}=p^{2n}$ given in Example 2.4. Observe that ${\Gamma\mathrm{L}}(3,p^{n/3})$ contains a Singer cycle, the cyclic subgroup of order $p^{n}-1$, and ${\mathrm{Aut}}(A_{p}(n,\theta))$ also has an automorphism of order $p^{n}-1$ given in Proposition 2.5 (ii). This is the key to prove $A_{p}(n,\theta)$ is a $3$-orbit group for $|\theta|=3$. We need the following lemma before proving the isomorphism.
Lemma 4.2.
Suppose that $V={\mathbb{F}}_{p}^{n}$. Let $G\leqslant{\mathrm{GL}}(V)\cong{\mathrm{GL}}(n,p)$ be the Singer cycle, the cyclic subgroup of ${\mathrm{GL}}(V)$ of order $p^{n}-1$. Then $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ is a direct sum of pairwise non-isomorphic irreducible ${\mathbb{F}}_{p}G$-submodules.
Proof.
Let ${\mathrm{F}}:={\mathbb{F}}_{p}$ and ${\mathrm{E}}:={\mathbb{F}}_{q}$. It is clear that $\Lambda^{2}_{{\mathrm{F}}}(V)$ is a direct sum of irreducible ${\mathrm{F}}G$-submodules by Maschke’s Theorem as $p\nmid|G|$. Note that $G\cong{\mathrm{GL}}(1,p^{n})$. Let $U$ be the natural ${\mathrm{E}}G$-module, and let $\phi$ be a generator of ${\mathrm{Gal}}({\mathrm{E}}/{\mathrm{F}})$. Then clearly that $U\cong V$ as ${\mathrm{F}}G$-modules. By Lemma 3.2, $\Lambda_{\mathrm{F}}^{2}(V)=\oplus_{i=1}^{\lfloor\frac{n}{2}\rfloor}B_{j}$ where
  1. (i)
    $B_{j}\cong U\otimes_{{\mathrm{E}}}U^{\phi^{j}}$ as ${\mathrm{F}}G$-modules for $1\leqslant j<\frac{n}{2}$;
  2. (ii)
    $2\cdot B_{n/2}\cong U\otimes_{{\mathrm{E}}}U^{\phi^{n/2}}$ as ${\mathrm{F}}G$-modules, when $n$ is even.
We claim that $B_{j}$’s are irreducible ${\mathrm{F}}G$-modules. Note that the kernel of $G$ on $B_{j}$ has order $\mathrm{gcd}(p^{j}+1,p^{n}-1)$. When $n$ is even, the induced permutation group of $G$ on $B_{n/2}$ is isomorphic to ${\mathbb{Z}}_{p^{n/2}-1}$. It follows that $B_{n/2}$ is an irreducible ${\mathbb{F}}_{p}G$-module as $\dim B_{n/2}=\frac{n}{2}$. Now we assume that $1\leqslant j<\frac{n}{2}$ and $n\geqslant 3$. When $n=6$ and $p=2$, it is not hard to see that both $B_{1}$ and $B_{2}$ are irreducible ${\mathrm{F}}G$-modules. When $(n,p)\neq(6,2)$ and $n\geqslant 3$, there exists a prime divisor $r$ of $p^{n}-1$ such that $r\mid p^{n}-1$ and $r\nmid p^{k}-1$ for any $k<n$ by Zsigmondy’s Theorem [33]. Then the order of the induced permutation group of $G$ on $B_{j}$ is divisible by $r$, and hence $B_{j}$ is an irreducible ${\mathbb{F}}_{p}G$-module as $\dim B_{j}=n$.
Since $\dim B_{n/2}\neq\dim B_{j}$ for $21\leqslant i\neq j<n/$ 2, we only need to show that $B_{i}\ncong B_{j}$ for $21\leqslant i\neq j<n/$ 2. If $B_{i}\cong B_{j}$ then $B_{i}^{\mathrm{E}}\cong B_{j}^{\mathrm{E}}$, which yields that $U\otimes_{\mathrm{E}}U^{\phi^{i}}\cong(U\otimes_{\mathrm{E}}U^{\phi^{j}})^{\phi^{k}}$ for some $k$. Let $g$ be a generator of $G$, and let $\lambda$ be the character of $g$ on $U$. Then $\lambda$ is a generator of ${\mathrm{E}}^{\times}$. Note that the character of $g$ on $U\otimes_{\mathrm{E}}U^{\phi^{i}}$ and $(U\otimes_{\mathrm{E}}U^{\phi^{i}})^{\phi^{k}}$ are $\lambda^{1+p^{i}}$ and $\lambda^{(1+p^{j})p^{k}}$, respectively. Then we have
$1+p^{i}\equiv p^{k}+p^{j+k}\pmod{p^{n}-1}.$
Let $\ell$ be the residue of $j+k$ modulo $n$. Then
$1+p^{i}\equiv p^{k}+p^{\ell}\pmod{p^{n}-1}.$
Note that $0\leqslant i,k,\ell<n$, we have
$-(p^{n}-1)<-(p^{i}-1)\leqslant p^{k}+p^{\ell}-p^{i}-1<p^{n}-1.$
Hence, the characters of $g$ on $U\otimes_{\mathrm{E}}U^{\phi^{i}}$ and $(U\otimes_{\mathrm{E}}U^{\phi^{i}})^{\phi^{k}}$ cannot be equal. Therefore, $B_{j}$’s are pairwise non-isomorphic irreducible summands of $\Lambda^{2}_{{\mathrm{F}}}(V)$. ∎
We remark that the character of ${\mathrm{F}}G$-module $V$ is equal to the character of $V^{\mathrm{E}}$ for any extension field ${\mathrm{E}}$ of ${\mathrm{F}}$. Let $U$ be an ${\mathrm{E}}G$-module such that $U\cong V$ as ${\mathrm{F}}G$-modules. Lemma 3.2 (i) deduces that
$\chi_{V}(g)=\mathrm{Tr}_{{\mathrm{E}}/{\mathrm{F}}}\chi_{U}(g)=\sum_{i=1}^{f}\chi_{U}(g)^{\phi^{i}},$
where $\chi_{V}$ and $\chi_{U}$ are characters of $V$ and $U$, respectively, and $\langle\phi\rangle={\mathrm{Gal}}({\mathrm{E}}/{\mathrm{F}})$ with $[{\mathrm{E}},{\mathrm{F}}]=f$. Now we are ready to get the 6-th row of Table 1.
Lemma 4.3.
Let $N:=A_{p}(n,\theta)$ be defined in Example 2.4 such that $3\mid n$, $|\theta|=3$ and $p$ is an odd prime. Then $N$ is a $3$-orbit group and ${\mathrm{Aut}}(N)\cong p^{n^{2}}{:}{\Gamma\mathrm{L}}(3,p^{n/3})$.
Proof.
Note that $N$ is a special $p$-group of exponent $p$. We assume that $N={\mathcal{H}}/W$ with ${\mathcal{H}}={\mathcal{H}}_{n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Let $V:={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})=N/{\mathbf{Z}}(N)$, and let $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)\cong{\mathrm{GL}}(n,p)$.
Recall that the automorphism $\xi\in{\mathrm{Aut}}(N)$, given in Proposition 2.5 (ii), has order $p^{n}-1$ and $\langle\xi\rangle$ acts faithfully on $V$. Hence, $\langle\xi\rangle$ induces a subgroup $H\leqslant X$ of order $p^{n}-1$ and $W^{H}=W$ by Lemma 2.10 (ii). Note that there exists $G\leqslant X$ such that $H<G$ and $G\cong{\mathrm{GL}}(3,q)$ with $q=p^{n/3}$. By Lemma 2.10 (i), 3.2 and 3.5, ${\mathbf{Z}}({\mathcal{H}})=A\oplus B$ as ${\mathbb{F}}_{p}G$-modules such that
  1. (i)
    $A\cong\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ as ${\mathbb{F}}_{p}G$-modules with $U$ the natural ${\mathbb{F}}_{q}G$-module; and
  2. (ii)
    $G$ acts transitively on non-zero vectors of ${\mathbf{Z}}({\mathcal{H}})/B\cong A$.
Thus, $N_{0}:={\mathcal{H}}/B$ is a $3$-orbit $p$-group with $|N_{0}|=|{\mathbf{Z}}(N_{0})|^{2}$ by Lemma 2.8 (iii) and 2.10 (ii). We claim that $W=B$, and then $N=N_{0}$ is a $3$-orbit $p$-group with ${\mathrm{Aut}}(N)^{V}\cong{\Gamma\mathrm{L}}(3,q)$ by Lemma 4.1.
Note that $B$ is an ${\mathbb{F}}_{p}H$-submodule of ${\mathbf{Z}}({\mathcal{H}})\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$. Hence, $B=W$ if and only if ${\mathbf{Z}}(N)$ and $A$ are isomorphic ${\mathbb{F}}_{p}H$-modules deduced by Lemma 4.2. Let $h\in H$ induced by $\xi$, and let $\chi_{{\mathbf{Z}}}$ and $\chi_{A}$ be characters of $h$ on ${\mathbf{Z}}(N)$ and $A$, respectively. It suffices to show that $\chi_{{\mathbf{Z}}}=\chi_{A}$. Let $\chi_{V}$ be the character of $h$ on $V$. By definition of $\xi$, we have
$\chi_{V}=\mathrm{Tr}_{{\mathbb{F}}_{p^{n}}/{\mathbb{F}}_{p}}(\lambda)\mbox{ and }\chi_{{\mathbf{Z}}}=\mathrm{Tr}_{{\mathbb{F}}_{p^{n}}/{\mathbb{F}}_{p}}(\lambda\cdot\lambda^{\theta}),$
where $\lambda$ is a generator of ${\mathbb{F}}_{p^{n}}^{\times}$. Note that $A\cong\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ as ${\mathbb{F}}_{p}H$-modules. Let $\mathcal{U}$ be the natural ${\mathbb{F}}_{p^{n}}H$-module. Then $\mathcal{U}\cong U$ as ${\mathbb{F}}_{q}H$-modules. Hence, Lemma 3.2 deduces that $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)\cong\mathcal{U}\otimes_{{\mathbb{F}}_{p^{n}}}\mathcal{U}^{\theta}$ as ${\mathbb{F}}_{q}H$-modules since $\theta$ is a generator of ${\mathrm{Gal}}({\mathbb{F}}_{p^{n}}/{\mathbb{F}}_{q})$. Then we have
$\chi_{A}=\mathrm{Tr}_{{\mathbb{F}}_{p^{n}}/{\mathbb{F}}_{p}}(\lambda\cdot\lambda^{\theta})=\chi_{{\mathbf{Z}}}.$
This yields that $W=B$. Therefore, $N$ is a $3$-orbit group and ${\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}\cong{\Gamma\mathrm{L}}(3,q)$ with $q=p^{n/3}$. ∎
4.2. ${\mathrm{Sp}}(2m,p^{n})\lhd G$
Now we treat the case that ${\mathrm{Sp}}(2m,q)\lhd G$. The following lemma shows that these groups are factor groups of $q^{1+2m}_{+}$.
Lemma 4.4.
Let $q:=p^{n}$ with $p$ an odd prime, and let $N$ be a special $p$-group of exponent $p$ with $N/{\mathbf{Z}}(N)\cong{\mathbb{Z}}_{p}^{2mn}$ for some integer $m$. The following statements hold.
  1. (i)
    If $|{\mathbf{Z}}(N)|=q$ and ${\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$ has a subgroup isomorphic to ${\mathrm{Sp}}(2m,q)$, then $N\cong q^{1+2m}_{+}$.
  2. (ii)
    If $|{\mathbf{Z}}(N)|\leqslant q$ and ${\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$ has a subgroup isomorphic to ${\mathrm{Sp}}(2m,q)$, then $N$ is a quotient of $q^{1+2m}_{+}$.
  3. (iii)
    ${\mathrm{Aut}}(q^{1+2m}_{+})\cong p^{2mn^{2}}{:}{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$.
Proof.
Let ${\mathcal{H}}:={\mathcal{H}}_{2mn}$, and let $V:={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$. Then $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)\cong{\mathrm{GL}}(2mn,p)$ and ${\mathbf{Z}}({\mathcal{H}})$ can be identify with $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ as ${\mathbb{F}}_{p}X$-modules by Lemma 2.10 (i). Note that $N={\mathcal{H}}/W$ for some $W<{\mathbf{Z}}(N)$. Then $W^{g}=W$ for $g\in G:={\mathrm{Aut}}(N)^{V}$ by Lemma 2.10 (ii). Suppose that $G_{0}\leqslant G$ isomorphic to ${\mathrm{Sp}}(2m,q)$ and $|{\mathbf{Z}}(N)|=|\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W|=q$. Note that $q^{1+2m}_{+}={\mathcal{H}}/W_{0}$ for some $W_{0}\leqslant{\mathbf{Z}}({\mathcal{H}})$ and $X_{W_{0}}$ contains a subgroup $G_{1}$ isomorphic to ${\mathrm{Sp}}(2m,q)$ by Lemma 2.10 (ii) and 2.19. Then there exists $x\in X$ such that $G_{1}^{x}=G_{0}$, and then $q^{1+2m}_{+}\cong{\mathcal{H}}/W_{0}^{x}$ with $G_{0}\leqslant X_{W_{0}^{x}}$. Thus, we may assume that $G_{1}=G_{0}$. It suffices to show that $W_{0}=W$. By Lemma 3.7 (iii), $S\subseteq W$ and $S\subseteq W_{0}$ where $S$ is an ${\mathbb{F}}_{p}G$-submodule of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ given in Lemma 3.7. Note that $\dim S=\dim\Lambda^{2}{{\mathbb{F}}_{q}}(V)-n$, it follows that $S=W=W_{0}$ as $|{\mathbf{Z}}({\mathcal{H}})/S|=|{\mathbf{Z}}(N)|=|{\mathbf{Z}}(q^{1+2m}_{+})|=p^{n}$. Thus, $N\cong q^{1+2m}_{+}$, as in part (i).
Suppose that $G_{0}\leqslant G$ isomorphic to ${\mathrm{Sp}}(2m,q)$ and $|{\mathbf{Z}}(N)|=|\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W|\leqslant q$. With the similar arguments in part (i), we obtain that $S\leqslant W$. Then $N$ is a quotient of ${\mathcal{H}}/S\cong q^{1+2m}_{+}$ as in part (ii).
Let $N:=q^{1+2m}_{+}$. If $q=p$, then ${\mathrm{Aut}}(N)\cong p^{2m}{:}\mathrm{CSp}(2m,q)$ given by Winter [31]. Suppose that $n\geqslant 2$. Then $G$ is non-solvable by Lemma 2.19. Note that $G$ acts transitively on non-zero vectors of both $V$ and $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ and $|V|\neq|\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W|$. Then $G$ has a normal subgroup isomorphic to ${\mathrm{Sp}}(2m^{\prime},p^{n^{\prime}})$ with $2m^{\prime}n^{\prime}=2mn$ by Theorem A. Since $G$ has a subgroup isomorphic to ${\mathrm{Sp}}(2m,q)$ Lemma 2.19, we have $n^{\prime}\leqslant n$. If $n^{\prime}<n$, then $|\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W|<q$ by Theorem A, a contradiction. Thus, $G$ has a normal subgroup isomorphic to ${\mathrm{Sp}}(2m,q)$. Therefore, $G={\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$ and we obtain part (iii) by Lemma 2.8 (i) and 2.11. ∎
We remark that the $1$-subspace stabilizer $G_{\alpha}$ of $G:={\mathrm{Sp}}(2m+2,q)$ is of form
$q^{1+2m}{:}({\mathrm{Sp}}(2m,q)\times{\mathrm{GL}}(1,q)),$
and $N:=\textbf{O}_{p}(G_{\alpha})\cong q^{1+2m}$ is a special $p$-group with $|{\mathbf{Z}}(N)|=q$ and ${\mathrm{Sp}}(2m,q)\leqslant{\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$ (see [30, Section 3.5.4]). These facts together with Lemma 4.4 (i) yield that $N\cong q^{1+2m}_{+}$, which is Theorem C (iii).
Corollary 4.5.
Let $G:={\mathrm{PSp}}(2m+2,q)$ acts naturally on $1$-spaces of ${\mathbb{F}}_{q}^{2m+2}$ where $q=p^{n}$ for odd prime $p$. Then $\textbf{O}_{p}(G_{\alpha})\cong q^{1+2m}_{+}$ for any stabilizer $G_{\alpha}$.
5. Subfield hyperplanes and quotients of $q^{1+2m}_{+}$
Let $q_{0}=p^{d}$ and $q=p^{n}$ with $d\mid n$. Note that ${\mathrm{Sp}}(2m,q)\leqslant{\mathrm{Sp}}(2mn/d,q_{0})$ for any positive integer $m$. Lemma 4.4 (ii) deduces that there exists $U<{\mathbf{Z}}(q^{1+2m}_{+})$ such that
$(q_{0})^{1+2mn/d}_{+}\cong q^{1+2m}_{+}/U.$
To describe the subgroup $U$, we introduce the following concept of subfield hyperplanes.
Definition 5.1.
Let $V={\mathbb{F}}_{q}^{+}$ with $q=p^{n}$ for some prime $p$, and let $x\in H$ be the linear transformation given by multiplying a generator $\lambda$ of ${\mathbb{F}}_{q}^{\times}$. For any divisor $d$ of $n$, we say an ${\mathbb{F}}_{p}$-subspace $U$ of $V$ is a subfield hyperplane with respect to subfield ${\mathbb{F}}_{p^{d}}$ if $\dim U=n-d$ and $U$ is invariant under $\langle x^{\ell}\rangle$ with $\ell=\frac{q-1}{p^{d}-1}$.
Note that $V={\mathbb{F}}_{p^{n}}^{+}$ is naturally an ${\mathbb{F}}_{p^{d}}$-space for any divisor $d$ of $n$. Hence, an ${\mathbb{F}}_{p}$-subspace $U$ of $V$ is a subfield hyperplane with respect to ${\mathbb{F}}_{p^{d}}$ if and only if $U$ is naturally a hyperplane of $V$ when viewing both $U$ and $V$ as ${\mathbb{F}}_{p^{d}}$-spaces. We provide the following lemma before studying $q^{1+2m}_{+}/U$.
Lemma 5.2.
Suppose that $H={\Gamma\mathrm{L}}(1,p^{n})$ acting naturally on $V={\mathbb{F}}_{p^{n}}^{+}$. For any divisor $d$ of $n$, the set of subfield hyperplane with respect to subfield ${\mathbb{F}}_{p^{d}}$ of $V$ forms an orbit of $H$.
In particular, $(H_{U})^{V/U}\cong{\Gamma\mathrm{L}}(1,p^{d})$ for any subfield hyperplane $U$ with respect to ${\mathbb{F}}_{p^{d}}$.
Proof.
Let $\Sigma$ be the set of subfield hyperplanes with respect to ${\mathbb{F}}_{p^{d}}$. Set $x\in H$ the operator admitted by multiplying a generator $\lambda\in{\mathbb{F}}_{p^{n}}^{\times}$, and set $y$ the operator admitted by the Frobenius automorphism in ${\mathrm{Aut}}({\mathbb{F}}_{p^{n}})$.
By definition, $U\in\Sigma$ if and only if $U$ is an ${\mathbb{F}}_{p}\langle x^{\ell}\rangle$-submodule of $V$ where $\ell=\frac{p^{n}-1}{p^{d}-1}$. Since $\langle x^{\ell}\rangle\lhd H$, $U^{h}$ is also an ${\mathbb{F}}_{p}\langle x^{\ell}\rangle$-submodule for any $h\in H$. Thus, $\Sigma$ is closed under the action of $H$.
Note that each $U\in\Sigma$ is a hyperplane of ${\mathbb{F}}_{p^{d}}^{n/d}$, and hence $|\Sigma|=\frac{p^{n}-1}{p^{d}-1}$. Since $\langle x\rangle_{U}=\langle x^{\ell}\rangle$, it follows that
$|U^{\langle x\rangle}|=\frac{|x|}{|x^{\ell}|}=\frac{p^{n}-1}{p^{d}-1}=|\Sigma|.$
Thus, $\langle x\rangle$ acts transitively on $\Sigma$, so does $H$.
Let $U\subseteq V={\mathbb{F}}_{p^{n}}^{+}$ consisting of elements $u$ such that
${\mathrm{Tr}}_{{\mathbb{F}}_{p^{n}}/{\mathbb{F}}_{p^{d}}}(u)=\sum_{k=1}^{n/d}u^{y^{dk}}=0.$
Clearly that $U$ is a hyperplane of ${\mathbb{F}}_{p^{d}}^{n/d}$ as ${\mathbb{F}}_{p^{d}}$-space. Thus, $U\in\Sigma$ and $U^{x^{\ell}}=U$. In particular, we have
${\mathrm{Tr}}_{{\mathbb{F}}_{p^{n}}/{\mathbb{F}}_{p^{d}}}(u^{y})={\mathrm{Tr}}_{{\mathbb{F}}_{p^{n}}/{\mathbb{F}}_{p^{d}}}(u)^{y}=0.$
Hence, $U^{y}=U$ and $U$ is an ${\mathbb{F}}_{p}\langle x^{\ell},y\rangle$-submodule of $V$. Since $\langle x^{\ell},y\rangle^{V/U}\cong{\Gamma\mathrm{L}}(1,p^{d})$, we complete the proof. ∎
Let $N=q^{1+2m}_{+}$. Then ${\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}\cong{\Gamma\mathrm{L}}(1,q)$ by Lemmas 2.19 and 4.4(iii). Hence, ${\mathbf{Z}}(N)$ can be naturally viewed as ${\mathbb{F}}_{q}^{+}$ by the action of ${\Gamma\mathrm{L}}(1,q)$. We are now ready to complete the proof of Theorem C.
Proof of Theorem C.
Let $N:=q^{1+2m}_{+}$, and let $N_{0}:=(q_{0})^{1+2mn/d}_{+}$. By Lemma 4.4 (iii), we have ${\mathrm{Aut}}(N_{0})^{N_{0}/{\mathbf{Z}}(N_{0})}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2mn/d,q_{0})$ which has a subgroup isomorphic to ${\mathrm{Sp}}(2m,q)$. Hence, $N_{0}\cong N/U_{0}$ for some $U_{0}<{\mathbf{Z}}(N)$ by Lemma 4.4 (i). Set $H:={\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}$. Then $H\cong{\Gamma\mathrm{L}}(1,q)$ by Lemma 2.19. Let $x$ be the element of order $q-1$ in $H$, and let $\delta$ be a preimage of $x$ in ${\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$. Then $\delta^{\ell}\in{\mathrm{Aut}}(N_{0})^{N_{0}/{\mathbf{Z}}(N_{0})}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2mn/d,q_{0})$ for $\ell=\frac{q-1}{q_{0}-1}$ (see [18, page 112]). Hence, $U_{0}$ is invariant under $\delta^{\ell}$, equivalently, $U_{0}^{x^{\ell}}=U_{0}$. Thus, $U_{0}<{\mathbf{Z}}(N)$ is a subfield hyperplane of with respect to ${\mathbb{F}}_{p^{d}}$ by definition.
Conversely, assume that $U<{\mathbf{Z}}(N)$ is a subfield hyperplane with respect to ${\mathbb{F}}_{p^{d}}$. Then there exists an $h\in H\cong{\Gamma\mathrm{L}}(1,q)$ such that $U^{h}=U_{0}$ by Lemma 5.2. Applying Lemma 2.12 by getting a preimage of $h$ in ${\mathrm{Aut}}({\mathcal{H}}_{2mn,p})$, we can obtain that
$q^{1+2m}_{+}/U\cong q^{1+2m}_{+}/U_{0}\cong(q_{0})^{1+2mn/d}_{+}.$
Notice that any subspace of codimension $1$ is a subfield hyperplane with respect to ${\mathbb{F}}_{p}$, then the “in particular” part immediately holds. The proof of part (i) is completed.
Additionally, parts (ii) and (iii) have been given in Lemma 2.15 and Corollary 4.5, respectively. ∎
Let $N=q^{1+2m}_{+}$ with $q=p^{n}$, and let $U<{\mathbf{Z}}(N)$. Suppose that $U$ contains a subfield hyperplane $U_{0}$ of ${\mathbf{Z}}(N)$ with respect to ${\mathbb{F}}_{q_{0}}$ with $q_{0}=p^{d}$,. By Theorem C (i), we have
$N/U\cong(N/U_{0})/(U_{0}/U)\cong(q_{0})^{1+2mn/d}_{+}/U_{1},$
where $U_{1}:=U_{0}/U$. Thus, we may focus on the case where $U$ contains no subfield hyperplanes, as in the following example.
Example 5.3.
For $q=p^{n}$ with odd prime $p$, let $N=q^{1+2m}_{+}/U$ such that
  1. (i)
    $U<{\mathbf{Z}}(q^{1+2m}_{+})\cong{\mathbb{F}}_{q}^{+}$ contains no subfield hyperplanes with respect to any subfield;
  2. (ii)
    ${\Gamma\mathrm{L}}(1,q)_{U}$ is transitive on ${\mathbf{Z}}(N)={\mathbf{Z}}(q^{1+2m}_{+})/U$.
The groups in the above example form the last row in Table 1.
Lemma 5.4.
Let $N=q^{1+2m}_{+}/U$ be defined in Example 5.3. Then $N$ is $3$-orbit group and
${\mathrm{Aut}}(N)\cong p^{2mn_{0}n}{:}({\mathrm{Sp}}(2m,q){:}{\Gamma\mathrm{L}}(1,q)_{U}),$
where $p^{n_{0}}=|{\mathbf{Z}}(N)|=q/|U|$.
Proof.
Let $V:=q^{1+2m}_{+}/{\mathbf{Z}}(q^{1+2m}_{+})$, and identify $N/{\mathbf{Z}}(N)$ with $V$. By Lemma 4.4 (iv), we have $G:={\mathrm{Aut}}(q^{1+2m}_{+})^{V}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$. Let $G_{0}$ be the normal subgroup of $G$ isomorphic to ${\mathrm{Sp}}(2m,q)$. Then $G_{0}$ acts trivially on ${\mathbf{Z}}(q^{1+2m}_{+})$ by Lemma 2.19. Thus, $G_{0}$ is a subgroup of ${\mathrm{Aut}}(N)^{V}$ as $U^{G_{0}}=U$. Hence, ${\mathrm{Aut}}(N)^{V}$ is transitive on non-zero vectors of $N/{\mathbf{Z}}(N)$.
Let $H:={\mathrm{Aut}}(q^{1+2m}_{+})^{{\mathbf{Z}}(q^{1+2m}_{+})}\cong G/G_{0}\cong{\Gamma\mathrm{L}}(1,q)$. Then $H_{U}\leqslant{\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}$. Since $H_{U}$ is transitive on non-zero vectors of ${\mathbf{Z}}(N)$, we have ${\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}$ is transitive on non-identity elements of ${\mathbf{Z}}(N)$. By Lemma 2.8 (iv), we obtain that $N$ is a $3$-orbit group.
Since $N$ is a $3$-orbit group, ${\mathrm{Aut}}(N)^{V}$ has a normal subgroup $G_{1}\cong{\mathrm{Sp}}(2m^{\prime},q_{0})$ with $q_{0}=p^{n^{\prime}}$ such that $m^{\prime}n^{\prime}=mn$ by Lemma 2.8 (iv) and Theorem A. Since $G_{0}\leqslant{\mathrm{Aut}}(N)^{V}$, we have $n^{\prime}\leqslant n$ and $G_{0}\leqslant G_{1}$. Suppose that $n^{\prime}>n$. By Lemma 4.4 (i), $N\cong(q_{0})^{1+2m^{\prime}}_{+}/U_{1}$ for some $U_{1}<{\mathbf{Z}}((q_{0})^{1+2m^{\prime}}_{+})$ where $q_{0}=p^{n^{\prime}}$. Then $(q_{0})^{1_{2}m^{\prime}}_{+}\cong q^{1+2m}_{+}/U_{2}$ with $U_{2}$ a subfield hyperplane of ${\mathbf{Z}}(q^{1+2m}_{+})$ with respect to ${\mathbb{F}}_{q_{0}}$ by Theorem C (i). Hence, we have
$q^{1+2m}_{+}/U\cong N\cong(q_{0})^{1+2m^{\prime}}_{+}/U_{1}\cong(q^{1+2m}_{+}/U_{2})/U_{1}.$
We may assume that the preimage of $U_{1}$ in ${\mathbf{Z}}(q^{1+2m}_{+})$ is $U$. Then $U_{2}<U$ is a subfield hyperplane, a contradiction. Thus, we have $n=n^{\prime}$ and $G_{0}=G_{1}\cong{\mathrm{Sp}}(2m,q)$. Therefore, ${\mathrm{Aut}}(N)^{V}\cong G_{0}{:}{\Gamma\mathrm{L}}(1,q)_{U}$, and then the lemma holds by Lemma 2.8 (i). ∎
Now we are in a position to prove Theorem B.
Proof of Theorem B.
Let $N$ be a finite $3$-orbit group. Note that the set of orders of elements of $N$ can only be $\{1,p,q\}$, $\{1,p,p^{2}\}$ and $\{1,p\}$ for some primes $p$ and $q$. Hence, $N$ is either a $\{p,q\}$-group or a $p$-group. If $N$ is a $\{p,q\}$-group, then Lemma 2.2 shows that $N$ is in the first row of Table 1. If $N$ is an abelian $p$-group, then $N$ is in the second row by Lemma 2.3. If $N$ is a non-abelian $2$-group, then $N$ is one of the groups in rows 3-5 by Theorem 2.6.
Assume that $N$ is a non-abelian $p$-group for odd prime $p$. Then $N$ is a special $p$-group of exponent $p$ by Lemma 2.7. If ${\mathrm{Aut}}(N)$ is solvable, then $N\cong q^{1+2}_{+}$ by Theorem 2.16 and is in the 7-th row of Table 1. If ${\mathrm{Aut}}(N)$ is non-solvable, then $G:={\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$ is non-solvable by Lemma 2.8 (i). Lemma 2.8 (iii) shows that $G$ acts transitively on non-identity elements of both $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$, and Lemma 2.8 (iv) shows that ${\mathbf{Z}}(N)$ is isomorphic to a quotient ${\mathbb{F}}_{p}G$-module of $\Lambda_{{\mathbb{F}}_{p}}^{2}(N/{\mathbf{Z}}(N))$. By Theorem A, one of the following cases hold.
  1. (1)
    $|{\mathbf{Z}}(N)|=1$;
  2. (2)
    $|{\mathbf{Z}}(N)|=|N/{\mathbf{Z}}(N)|$;
  3. (3)
    $G$ has a normal subgroup $G_{0}\cong{\mathrm{Sp}}(2m,p^{n})$, and $|{\mathbf{Z}}(N)|\leqslant p^{n}$.
If (1) holds, then $N$ is an extraspecial $p$-group of exponent $p$, which is in the 7-th row. If (2) holds, then $N$ is in the 6-th row by Lemmas 4.1 and 4.3. Assume that (3) holds. Then $N$ is a quotient of $q^{1+2m}_{+}/U$ for some $U\leqslant{\mathbf{Z}}(q^{1+2m}_{+})$ by Lemma 4.4 (ii), and $q^{1+2m}_{+}$ is listed in the 7-th row. If $U\neq 1$ and contains some subfield hyperplane $U_{0}$, then $N$ is a quotient of $q^{1+2m}_{+}/U_{0}\cong(q_{0})^{(1+2mn_{0})}_{+}$ by Theorem C (i). Thus, we may assume that $U$ contains no subfield hyperplanes. Then, Lemma 5.4 deduces the last row of Table 1, and we complete the proof. ∎
We now prove Corollary 1.1.
Proof of Corollary 1.1.
Let $N_{1}:=N/U$, and let $V:=N/{\mathbf{Z}}(N)=N_{1}/{\mathbf{Z}}(N_{1})$. Then ${\mathrm{Aut}}(N)^{V}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$ by Lemma 4.4 (iii). Since ${\mathrm{Sp}}(2m,q)\leqslant{\mathrm{Aut}}(N)^{V}$ acts trivially on ${\mathbf{Z}}(N)$, it follows that ${\mathrm{Sp}}(2m,q)\leqslant{\mathrm{Aut}}(N_{1})^{V}$ and ${\mathrm{Aut}}(N_{1})^{V}$ acts transitively on non-zero vectors of $V$. Thus, $N_{1}$ is $3$-orbit if and only if ${\mathrm{Aut}}(N_{1})^{{\mathbf{Z}}(N_{1})}$ is transitive on non-identity elements of ${\mathbf{Z}}(N_{1})={\mathbf{Z}}(N)/U$ deduced by Lemma 2.8 (iii). It suffices to prove that $(H_{U})^{{\mathbf{Z}}(N_{1})}={\mathrm{Aut}}(N_{1})^{{\mathbf{Z}}(N_{1})}$.
Let $U_{0}$ be a subfield hyperplane with respect to subfield ${\mathbb{F}}_{p^{d}}$ for some $d\mid n$ such that $U_{0}\leqslant U$ and $U_{1}:=U/U_{0}$ has no subfield hyperplanes of ${\mathbf{Z}}(N)/U_{0}\cong{\mathbb{F}}_{p^{d}}^{+}$. By Theorem C (i) and Lemma 4.4 (iii), we have
$N_{0}:=N/U_{0}\cong(q_{0})^{1+2mn/d}_{+}\mbox{ and }H_{0}:={\mathrm{Aut}}(N_{0})^{{\mathbf{Z}}(N)/U_{0}}\cong{\Gamma\mathrm{L}}(1,q_{0}).$
Note that $H_{U_{0}}^{{\mathbf{Z}}(N)/U_{0}}=H_{0}\cong{\Gamma\mathrm{L}}(1,q_{0})$ by Lemma 5.2. Hence, we have
$(H_{U})^{{\mathbf{Z}}(N_{1})}=(H_{U})^{{\mathbf{Z}}(N)/U}=(H_{0})_{U_{1}}^{{\mathbf{Z}}(N_{0})/U_{1}}=(H_{0})_{U_{1}}^{{\mathbf{Z}}(N_{1})}.$
Since $U_{1}$ has no subfield hyperplanes, Lemma 5.4 deduces that
${\mathrm{Aut}}(N_{1})^{{\mathbf{Z}}(N_{1})}=((H_{0})_{U_{1}})^{{\mathbf{Z}}(N_{1})}=(H_{U})^{{\mathbf{Z}}(N_{1})}.$
Therefore, $N/U$ is $3$-orbit if and only if $H_{U}$ is transitive on ${\mathbf{Z}}(N)/U$. ∎
At last of this paper, we construct two families of groups in the 8-th row of Table 1.
Let $V={\mathbb{F}}_{p^{n}}^{+}$, and let $1\neq U<V$. Note that $U$ is a subfield hyperplane when $n=2$. For $n=3$, $U$ contains no subfield hyperplanes if and only if $\dim U=1$. If $\dim U=1$, then ${\mathrm{GL}}(1,p^{3})_{U}\cong{\mathbb{Z}}_{p-1}$, and hence $|{\Gamma\mathrm{L}}(1,p^{3})_{U}|\leqslant 3(p-1)$. Notice that $|V/U|-1=p^{2}-1>3(p-1)$ in this case. Therefore, smallest groups $q^{1+2m}_{+}/U$ in Example 5.3 satisfies $q\geqslant 3^{4}$.
When $V={\mathbb{F}}_{3^{4}}^{+}$, let $\lambda$ be a generator of ${\mathbb{F}}_{3^{4}}^{\times}$ with minimal polynomial $\lambda^{4}-\lambda^{3}-1=0$. Assume that $x$ is the operator of multiplying $\lambda$. Then $V$ is a direct sum of the following two subspaces:
$U_{1}:=\langle 1+\lambda^{2},2\lambda+\lambda^{2}+2\lambda^{3}\rangle\mbox{ and }U_{2}:=\langle 2+\lambda,1+\lambda+\lambda^{2}\rangle.$
We remark that each of $U_{1}$ and $U_{2}$ contains no subfield hyperplanes, and $H_{U_{i}}\cong{\mathbb{Z}}_{8}$ acts transitively on non-zero vectors of $V/U_{i}$ for $i=1,2$ where $H={\Gamma\mathrm{L}}(1,3^{4})$. Hence, $U_{1}$ and $U_{2}$ admit two smallest $3$-orbit groups (of order $3^{10}$) in the 8-th row of Table 1, which are $3^{1+2}_{+}/U_{1}$ and $3^{1+2}_{+}/U_{2}$, respectively. Notice that there exists $h\in N_{H}(L)$ such that $U_{1}^{h}=U_{2}$, the above two groups are isomorphic. We present a generating relation for the above group blow:
$N=\langle$ $x_{i},y_{j},z_{k}\mbox{ for $1\leqslant i,j,k\leqslant 4$ and $1\leqslant k\leqslant 4$}\mid x_{i}^{3}=y_{i}^{3}=z_{i}^{3}=1\mbox{ and }$
$[x_{i},x_{j}]=[y_{i},y_{j}]=[z_{i},x_{j}]=[z_{i},y_{j}]=1\mbox{ for each $1\leqslant i,j\leqslant 4$;}$
$[x_{1},y_{1}]=z_{1},\ [x_{1},y_{2}]=[x_{2},y_{1}]=z_{2},\ [x_{1},y_{3}]=[x_{2},y_{2}]=[x_{3},y_{1}]=z_{3},$
$[x_{1},y_{4}]=[x_{2},y_{3}]=[x_{3},y_{2}]=[x_{4},y_{1}]=z_{4},\ [x_{2},y_{4}]=[x_{3},y_{3}]=[x_{4},y_{2}]=z_{1}z_{4},$
$[x_{3},y_{4}]=[x_{4},y_{3}]=z_{1}z_{2}z_{4},\ [x_{4},y_{4}]=z_{1}z_{2}z_{3}z_{4},\ z_{1}z_{3}=z_{2}^{2}z_{3}z_{4}^{2}=1\rangle.$
We thank to Professor O’Brien, who told us that the TameGenus package in Magma  provides some functions dealing with $p$-groups having nilpotency class $2$, exponent $p$ and center of order $p^{2}$. The readers can use the TGAutomorphismGroup command in Magma to check that ${\mathrm{Aut}}(N)\cong 3^{2\times 8}{:}({\mathrm{Sp}}(2,3^{4}){:}{\mathbb{Z}}_{8})$.
We now construct another family of groups in the 8-th row of Table 1 by obtaining a family of desired subspaces $U$ of ${\mathbb{F}}_{q}^{+}$.
Example 5.5.
Let $p$ and $r$ be distinct odd primes such that $r\nmid p-1$, and let $q=p^{n}$ with $n=p^{r}-1$. Set $V:={\mathbb{F}}_{q}^{+}$, and let $H={\Gamma\mathrm{L}}(1,q)$ acting naturally on $V$. Let $y\in H$ admitted by the Frobenius automorphism in ${\mathrm{Aut}}({\mathbb{F}}_{q})$. Then there exists an $r$-dimensional faithful ${\mathbb{F}}_{p}\langle y\rangle$-submodule $R$ of $V$. Let $U$ be an ${\mathbb{F}}_{p}L$-complement of $R$ in $V$.
With definitions in the above example, we observe that $|V/U|-1=|R|-1=p^{r}-1=|y|$. Hence, $\langle y\rangle$ acts transitively on non-zero vectors of $V/U$, and then $q^{1+2m}_{+}/U$ is a $3$-orbit group for any positive integer $m$ by Corollary 1.1. Since $\dim R=r$ is not a divisor of $\dim V=p^{r}-1$, $q^{1+2m}_{+}/U$ lies in 8-th row of Table 1. Remark that the minimum integer $n$ satisfying the conditions in Example 5.5 is $n:=p^{r}-1=5^{3}-1=124$. Thereofore, the smallest $3$-orbit group admitted by Example 5.5 has order $5^{251}$ with center of order $5^{3}$.